### v1

Abstract
Equiangular tight frames (ETFs) are optimal packings in projective space which also yield useful decompositions of signals. Paley ETFs are constructed using number theory. In this article, we present the doubly homogeneous automorphism groups of the Paley ETFs of prime order. We also prove some properties of doubly homogeneous frames, namely that they are always ETFs and that their short circuits—i.e., dependent subsets of minimum size—form a balanced incomplete block design. We additionally characterize all $k$-homogeneous ETFs for $3\leq k\leq 5$. Finally, we revisit David Larson’s AMS Memoirs Frames, Bases, and Group Representations [HL00] coauthored with Deguang Han and Wandering Vectors for Unitary Systems and Orthogonal Wavelets [DL98] coauthored with Xingde Dai with a modern eye and focus on finite-dimensional Hilbert spaces. Keywords: equiangular tight frames, ETFs, $k$-homogeneous, $k$-transitive, Paley difference sets, $t$-designs, matroid circuits
1 Introduction
Frames are generalizations of orthonormal bases that have found wide applications in signal and image processing, compressed sensing, phase retrieval, and even characterizing emergent phenomenon in neural networks [CK13, KC07, FR13, PHD20]. Frames were originally introduced in [DS52] as a tool to loosen the definition of Fourier series. Frames over finite dimensional Hilbert spaces (see, e.g., [Wal18]), in particular so-called equiangular tight frames (ETFs) are intimately related to open problems in quantum information theory (like Zauner’s conjecture [Zau99, Zau11]) and combinatorial design theory (like the Gillespie conjecture [Gil18]). In this paper, we prove algebraic and combinatorial properties of an infinite class of ETFs (Paley ETFs). We also completely characterize all ETFs which have $3$-, $4$-, or $5$-homogeneous automorphism groups. Finally, we re-examine two of Dave Larson’s fundamental manuscripts from the turn of the millennium [HL00, DL98] in the context of frames in finite dimensions.
The author first learned about frames in 2002 in Dave Larson’s Research Experiences for Undergraduates (REU) on Frames and Wavelets, which ran for many years and inspired numerous young mathematicians. She has also used his book Frames for undergraduates [HKLW07]—cowritten with Deguang Han, Keri Kornelson, and Eric Weber, which came out of the experience running the REU—when working with students from high school to master’s level. Almost a quarter of a century later, the REU run by Dave Larson still has a large impact on her research program (and the research of many other alumni of the REU).
The basics of equiangular tight frames are covered in Section 1.1, followed by sections on their symmetries Section 1.2 and linear dependencies Section 1.3. In Definition 1.22 in Section 1.3, a new combinatorial concept is introduced: the bender, which consists of the short circuits of an ETF, which are the dependent subsets of minimum size. Section 2 contains results concerning Paley ETFs, with the main result being Theorem 2.8, in which the automorphism group of an infinite class of Paley ETFs is completely characterized. Section 3 concerns $k$-homogeneous ETFs, including Theorem 3.6 which explicitly lists every $3$-, $4$-, and $5$-homogenous ETF. Finally, some key results in Dave Larson’s Memoirs [HL00, DL98] are contextualized within the framework of modern research on finite frames in Section 4. This section also includes a list of some questions and conjectures.
1.1 Equiangular Tight Frames
For any separable Hilbert space $\mathcal{H}$ and any orthonormal basis $\Phi=(\varphi_{j})_{j\in J}$, Parseval’s equality yields that for all $x\in\mathcal{H}$
$\left\|x\right\|^{2}=\sum_{j\in J}\left|\langle x,\varphi_{j}\rangle\right|^{2}.$
What might be surprising is that there are collections of vectors $\Phi$ which are not orthonormal bases which also satisfy that equality exactly (Parseval frames), in a scaled version (tight frames), or in a loosened version (frames). In infinite dimensions, even being a frame requires significant structure, and there is an immense amount of research about such vectors (see, e.g., [Chr16]).
Definition 1.1.
For a separable Hilbert space $\mathcal{H}$ and $\Phi=(\varphi_{j})_{j\in J}$ in $\mathcal{H}$, $\Phi$ is a tight frame if there exists $A>0$ such that for all $x\in\mathcal{H}$
$A\left\|x\right\|^{2}=\sum_{j\in J}\left|\langle x,\varphi_{j}\rangle\right|^{2}.$
$A$ is called the frame bound.
In finite dimensions, any spanning set is a frame, and one often asks for additional structure, like being an equiangular tight frame. We will use the notation $\mathbb{F}$ to mean $\mathbb{R}$ or $\mathbb{C}$ throughout the paper.
Definition 1.2.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. We call $\Phi$ equiangular if there exist $0\leq\alpha\leq\beta$ such that
$\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(\beta-\alpha)\delta_{j,k}+\alpha.$
If additionally, $\Phi$ is a tight frame, then we call $\Phi$ an equiangular tight frame (ETF) or a $(d,n)$-ETF.
In $\mathbb{F}^{d}$, it is common to employ the abuse of notation using $\Phi$ to both mean the frame itself and the short fat matrix with columns the frame vectors. With that notation, $\Phi$ is a tight frame precisely when $\Phi\Phi^{\ast}=AI$. Equiangular tight frames are optimally robust to noise and up to two erasures [GKK01, SH03, HP04], are important in quantum information theory [Zau99, Zau11], and provide optimal projective codes [JKM19]. As mathematical objects which interact with a variety of fields, there are a number of conjectures about their existence or the lack thereof: Zauner’s conjecture [Zau99, Zau11], the Gillespie conjecture [Gil18], the $(d,2d)$ conjecture [FI23, IJM24], and the Fickus-Jasper conjecture [FJ19].
For many choices of $d$, $n$, $\mathbb{F}$, there does not exist an ETF. (See, e.g, [FM16].) When $n=d$, the ETFs are simply equal-norm orthogonal bases. When $n=d+1$, they span lines which go through vertices of a regular simplex.
Definition 1.3.
If $\Phi$ is an ETF of $d+1$ vectors in $\mathbb{F}^{d}$, we call it a simplex ETF. If further, the inner product of any two distinct vectors in a simplex ETF is negative (necessarily the same negative number since they are equiangular), then we call it a regular simplex ETF.
In a certain sense, every simplex ETF is equivalent to a regular simplex ETF. Let $\operatorname{U}(d)$ denote the $d\times d$ orthogonal matrices when $\mathbb{F}=\mathbb{R}$ and $d\times d$ unitary matrices when $\mathbb{F}=\mathbb{C}$. Then there is always a $U\in\operatorname{U}(d)$ and a choice of scalars $\eta_{j}\in\operatorname{U}(1)$ such that
$\left(U\varphi_{j}\eta_{j}\right)_{j=1}^{d+1}$
have real entries and point to vertices of a regular $d$-simplex centered at the origin. In general, one may ask if two sets of vectors are fundamentally “the same”, i.e., span the same set of lines up to a global change of basis. Fortunately, there is a relatively simple test.
Definition 1.4.
Assume that $\Phi=(\varphi_{j})_{j=1}^{n},\Psi=(\psi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. We say that the vectors are switching equivalent (or projectively unitarily equivalent or isometrically isomorphic) if there exists a $U\in\operatorname{U}(d)$ and a choice of scalars $\eta_{j}\in\operatorname{U}(1)$ such that
$\psi_{j}=U\varphi_{j}\eta_{j},\quad j=1,2,\ldots,n.$
The test for switching equivalence has been rediscovered in different forms over many years (e.g., [Sei76, GP77, Bre90, AFF11, CW16])
Proposition 1.5.
Let $\Phi=(\varphi_{j})_{j=1}^{n},\Psi=(\psi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ be ETFs. They are switching equivalent if and only if for $j\neq k\neq\ell\neq j$
$\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle=\left\langle\psi_{j},\psi_{k}\right\rangle\left\langle\psi_{k},\psi_{\ell}\right\rangle\left\langle\psi_{\ell},\psi_{j}\right\rangle.$
We will use the notation $\operatorname*{TP}(j,k,\ell)=\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle$, where $\Phi$ will be clear from context.
For $d\neq n$, there exists a $(d,n)$-ETF over $\mathbb{F}$ if and only if there exists a $(n-d,n)$-ETF over $\mathbb{F}$. Such frames are called Naimark complements of each other. Naimark complements of Parseval frames (not in general equiangular and allowing infinite dimensional Hilbert spaces) were first studied by Han and Larson [HL00], where they were called complementary frames. It was later realized that these results followed from Naimark’s dilation theorem concerning the lifting of positive operator valued measures [Neu43].
Proposition 1.6.
Let $\Phi$ be a $(d,n)$-ETF over $\mathbb{F}$ with frame bound $A$. Then any $\Psi\in\mathbb{F}^{(n-d)\times n}$ that satisfies
$\Psi^{\ast}\Psi=AI-\Phi^{\ast}\Phi$
is an $(n-d,n)$-ETF over $\mathbb{F}$ with frame bound $A$ called a Naimark complement of $\Phi$.
Note that the inner products of distinct vectors of a Naimark complement are simply the negatives of the corresponding vectors in the first frame. Thus, by Proposition 1.5, characterizing switching equivalence of an ETF is equivalent to solving the problem for Naimark complements. Hence, it suffices to study ETFs for which $n\geq 2d$.
Given the historical connection of frames to Fourier bases [DS52], it is perhaps not surprising that Fourier-related concepts arise in constructions of various “nice” frames, like frames in $L^{2}(\mu)$ for $\mu$ a fractal measure (see, e.g., [DHSW11]) or Gabor frames for $L^{2}(\mathbb{R}^{d})$ (see, e.g., [Grö01]). Discrete Fourier transform matrices may also be used to define “nice” frames in finite dimensions.
Definition 1.7.
[…]
Note that this construction via Kronecker products yields a natural way to view the rows of the Fourier matrix as being indexed lexicographically by the elements of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$. By the fundamental theorem of finitely generated abelian groups, the character table of any finite abelian group is (up to possible permutation of entries) a Fourier matrix. Any $n\times n$ Fourier matrix is a unitary matrix scaled by $\sqrt{n}$; thus, any $d$-row all-column (i.e., $n$-column) submatrix $\Phi$ of an $n\times n$ Fourier matrix satisfies $\Phi\Phi^{\ast}=nI_{d}$. So, such submatrices always yield tight frames. We can completely characterize when such submatrices yield ETFs using combinatorial design theory (see, e.g., [CD07]).
Definition 1.8.
Let $G$ be a finite abelian group of size $v$ written additively. A $(v,k,\lambda)$-difference set is a subset $X\subseteq G$ of size $k$ such that the multiset of differences contains each non-identity element exactly $\lambda$ times.
Example 1.9.
Consider $\{1,2,4\}\subset\mathbb{Z}_{7}$. Then the difference table is
$\begin{array}[]{c||c|c|c}-&1&2&4\\ \hline\cr\hline\cr 1&0&1&3\\ \hline\cr 2&6&0&2\\ \hline\cr 4&4&5&0\end{array}$
That is, the multiset of differences contains each non-identity element of $\mathbb{Z}_{7}$ exactly once. Thus $\{1,2,4\}$ is a $(7,3,1)$-difference set.
In any abelian group of size $v$, any subset of size $1$, $v-1$, or $v$ trivially forms a difference set. The difference set in Example 1.9 is the smallest non-trivial difference set. In particular, not all abelian groups have a nontrivial difference set. Examples of the construction of ETFs using difference sets appeared in [SH03, DF07, GR09], while the following theorem which completely characterizes the construction may be found in [XZG05].
Theorem 1.10.
Let $G$ be a finite abelian group with Fourier matrix $F$. Let $\Phi$ be a all-column submatrix of $F$, with rows indexed by $X$. Then $\Phi$ is an ETF if and only if $X$ is a difference set in $G$.
1.2 The Geometric Structure of ETFs
One way to further understand ETFs is by characterizing their geometric and combinatorial properties via their symmetries and linear dependencies. We introduce the former concept in this section and the latter in the next section.
Definition 1.11.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a frame for $\mathbb{F}^{d}$. We call $U\in\operatorname{U}(d)$ a vector automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}=\varphi_{\sigma(j)}$ for $j=1,\ldots n$. We call $U\in\operatorname{U}(d)$ a line automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}\varphi_{j}^{\ast}U^{\ast}=\varphi_{\sigma(j)}\varphi_{\sigma(j)}^{\ast}$ for $j=1,\ldots n$. The group of vector automorphisms is the vector automorphism group $G$ of $\Phi$ and the quotient group of the line automorphisms modulo $\operatorname{U}(1)$ is the line automorphism group $[G]$ of $\Phi$.
By abuse of notation, we will work with elements of $\operatorname{U}(d)$ rather than $\operatorname{PU}(d)$ when analyzing the line automorphisms. In particular, $\operatorname{U}(1)$ plays no role in what permutation of the lines is induced.
Example 1.12.
Let $G$ be a finite abelian group with $n\times n$ Fourier matrix $F$. Let $\Phi$ be a $d$-row, all-column submatrix of $F$, which is necessarily a tight frame. Further, define $M_{j}$ for each frame index $j$ as the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. Then $\{M_{j}\}_{j}$ is a group in $\operatorname{U}(d)$ which is a representation of $G$. Also, $F$ is the orbit of the all-ones vector $\mathbbm{1}$ under $\{M_{j}\}_{j}$, where $M_{j}\mathbbm{1}=\varphi_{j}$.
We will often ask when the automorphism groups have extra structure.
Definition 1.13.
An action of a group $G$ on a set $X=\left\{\,x_{i}\,\,\middle|\,\,i\in\{1,\ldots,n\}\,\right\}$ is $k$-transitive if for each ordered $k$-tuple with no repeated elements $(x_{i_{1}},\ldots,x_{i_{k}})$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as ordered tuples)
$(g\cdot x_{i_{1}},\,g\cdot x_{i_{2}},\,\ldots,\,g\cdot x_{i_{k}})=(x_{\sigma(i_{1})},\,x_{\sigma(i_{2})},\,\ldots,\,x_{\sigma(i_{k})}).$
The action is $k$-homogeneous if for each set of $k$ elements of $X$ with no repeated elements $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as sets)
$\{g\cdot x_{i_{1}},\,g\cdot x_{i_{2}},\,\ldots,\,g\cdot x_{i_{k}}\}=\{x_{\sigma(i_{1})},\,x_{\sigma(i_{2})},\,\ldots,\,x_{\sigma(i_{k})}\}.$
A frame is $k$-transitive (respectively, $k$-homogeneous) if its line automorphism group is $k$-transitive (respectively, $k$-homogeneous).
Clearly, $k$-transitivity implies both $(k-1)$-transitivity and $k$-homogeneity. However, $k$-homogeneity does not nest. Consider a set of size $k$ acted on by the trivial group. Then the group action is $k$-homogeneous but not $j$-homogeneous for any $1\leq j<k$.
We make a note of a result from [CW18] (cf. Proposition 1.5) that we will leverage to prove a few results and which was also the basis of the code provided by Joey Iverson that was used to test some ideas.
Theorem 1.14.
[CW18] Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. Then there is a $g\in G$ inducing a permutation $\sigma\in S_{n}$ of the vectors if and only if for all $j\neq k$
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\left\langle\varphi_{\sigma(j)},\varphi_{\sigma(k)}\right\rangle.$
Similarly, there is a $g\in[G]$ inducing a permutation $\sigma\in S_{n}$ of the lines if and only if for all $j\neq k\neq l\neq j$
$\operatorname*{TP}(j,k,\ell)=\operatorname*{TP}(\sigma(j),\sigma(k),\sigma(\ell)).$
Since this result relies solely on the inner products, one can see that both the vector and the line automorphism groups of an ETF and a Naimark complement of an ETF are the same.
1.3 The Combinatorial Structure of ETFs
Understanding the linear dependencies of vectors in ETFs rose to prominence in the search for so-called deterministic restricted isometry property (RIP) measurements in compressed sensing (see, e.g., [BFMW13]). However, it was realized that many ETFs have in some sense worst possible linear dependence properties [FMT12].
Definition 1.15.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. The spark (or girth) of $\Phi$, $\operatorname{spark}(\Phi)$, is the size of the smallest subset of linearly dependent vectors in $\Phi$.
“Spark” is the term from compressed sensing literature [DE03], while “girth” is a more classical term from matroid theory (see, e.g.,[Oxl11]). RIP concerns subsets up to a certain size of columns of a matrix having condition number close to $1$; so, small subsets of columns being linearly dependent means that the matrix is far from being RIP.
Example 1.16.
Consider an orthonormal basis $(u_{j})_{j=1}^{d}$ for $\mathbb{F}^{d}$. Then
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&\sum_{j=1}^{d}u_{j}\end{pmatrix}=d+1$,
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&u_{1}\end{pmatrix}=2$, and
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&0\end{pmatrix}=1$,
where the singleton set consisting of the zero vector is counted as a linearly dependent set of size $1$, while the rank of each of the vector collections above is $d$.
Trivially, the largest that the spark of $n$ vectors in $\mathbb{F}^{d}$ may be is $d+1$, which happens when any subset of $d$ vectors is linear independent. When this happens, we say the vectors are full spark. Full spark frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ form an open dense set in the space of frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ [Cah09, ACM12, CMS17]. However, the same cannot be said when restricting to ETFs. Applying a now classical result from the compressed sensing literature [DE03] to ETFs, we have the following result.
Proposition 1.17.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a $(d,n)$-ETF. Then
$\operatorname{spark}(\Phi)\geq 1+\sqrt{\frac{d(n-1)}{n-d}}.$
The infinite class of ETFs constructed in [FMT12] all saturate this worst-case bound for spark. Inspired by this, the author of this paper and coauthors characterized all ETFs which saturate this bound [FJKM18].
Theorem 1.18.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a $(d,n)$-ETF. Then
$\operatorname{spark}(\Phi)=1+\sqrt{\frac{d(n-1)}{n-d}}$
if and only if $\Phi$ contains a subset of vectors which form a simplex ETF for their span.
This lead to the following definition [FJKM18].
Definition 1.19.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. The binder of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which form simplex ETFs for their span.
If the binder is non-empty, then each of the elements of the binder (which are subsets of $\Phi$) have the same size. In certain cases (e.g., Gabor Steiner frames [BK19]), the binder forms a type of combinatorial design call a BIBD. For more on combinatorial designs see, e.g., [CD07].
Definition 1.20.
A $t$-$(v,k,\lambda)$-design or block design is an ordered pair $(V,\mathcal{B})$, where $V$ is a set of size $v$ and $\mathcal{B}$ is a collection of subsets of $V$ called blocks such that each block $\beta\in\mathcal{B}$ has the same size $k$ and any subset of $t$ distinct elements in $V$ is contained in precisely $\lambda$ blocks. When $t=2$, the design is called a balanced incomplete block design (BIBD).
Note that $t$-designs are also $s$-designs for any $1\leq s\leq t$.
Example 1.21.
Consider the following subsets of $\mathbb{Z}_{7}$ which result from shifting the difference set in Example 1.9 by each element of the group:
$\left\{\{1,2,4\},\,\{2,3,5\},\,\{3,4,6\},\,\{0,4,5\},\,\{1,5,6\},\,\{0,2,6\},\,\{0,1,3\}\right\}.$
Note that $\{1,2\}$ is only in the first block, $\{1,3\}$ only in the seventh, $\{1,4\}$ only in the first, and so on. This is a $2$-$(7,3,1)$-design a.k.a. a $(7,3,1)$-BIBD.
The binder, however, of most ETFs is empty. This does not mean that the ETF is full spark, only that it does not have worst case spark. So, one may ask to study the properties of the smallest dependent sets. This leads us to introduce a new definition.
Definition 1.22.
Let $\Phi$ be an ETF with spark $s$. The bender111“Bender” is in honor of the smack-talking robot on Futurama, a show which is filled with math jokes. of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which are size $s$ and linearly dependent. The elements of the bender are called short circuits.
In matroid theory, minimally dependent subsets of vectors, i.e., those for which any proper subset is independent, are called circuits. The size of the smallest circuit(s) is the girth, i.e., the spark. Hence the name “short circuit.”
If $\Phi$ has a non-trivial binder, then the bender is the binder and the short circuits are all simplex ETFs for their span. On the other extreme, if $\Phi$ is full spark $d+1$, then the bender is the set of all subsets of size $d+1$ of the vectors.
It was noted in [FL23] that the binder of an ETF with a doubly transitive line automorphism group is a BIBD. However, folklore in combinatorial design theory (see, e.g., [CD07, Remark 4.29]) means a much more general result holds.
Proposition 1.23.
Let $\Phi$ be an ETF with $k$-homogeneous line automorphism group. Then the bender forms a $k$-design.
Proof.
Clearly, acting on a short circuit with an element of the line automorphism group results in a short circuit since unitaries do not change linear dependency. For any two different sets $X_{1}$ and $X_{2}$ of $k$ frame vectors, there is a least one element of the line automorphism group mapping $X_{1}$ to $X_{2}$, meaning that the size of the set of short circuits containing $X_{1}$ is the same as the size of the set of short circuits containing $X_{2}$. Thus, the bender is a $k$-design. ∎
2 Paley ETFs
We will apply the tools from Sections 1.2 and 1.3 to combinatorially and geometrically characterize one infinite class of ETFs.
We begin by reviewing some basic facts about finite fields. If $q=p^{s}$ for some prime $p$ and $\mathbb{F}_{q}$ is the finite field of size $q$, then $\mathbb{F}_{q}^{\times}=\mathbb{F}_{q}\backslash\{0\}$ is a cyclic group under multiplication. We denote the quadratic residues and non-residues as
$\operatorname{QR}(\mathbb{F}_{q})$ =$=\left\{\,a^{2}\,\,\middle|\,\,a\in\mathbb{F}_{q}^{\times}\,\right\}$
$\operatorname{NR}(\mathbb{F}_{q})$ $=\mathbb{F}_{q}^{\times}\backslash\operatorname{QR}.$
Thus, $\mathbb{F}_{q}=\operatorname{QR}(\mathbb{F}_{q})\cup\operatorname{NR}(\mathbb{F}_{q})\cup\{0\}$. When the underlying field is clear from context, we will often simply write $\operatorname{QR}$ and $\operatorname{NR}$. Note that $\operatorname{QR}$ is thus a subgroup of the multiplicative cyclic group $\mathbb{F}_{q}^{\times}$ of index $2$, meaning it has size $(q-1)/2$ and is also cyclic. When $q\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$; that is, the non-quadratic residues are precisely the negations of the quadratic residues. This fact implies the following (see, e.g. [CD07]).
Theorem 2.1.
When $q\equiv 3\bmod 4$, $\operatorname{QR}(\mathbb{F}_{q})$ forms a difference set in the additive group of $\mathbb{F}_{q}$.
Example 2.2.
We have that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ since in $\mathbb{F}_{7}$
$1^{1}=1,\quad 2^{2}=4,\quad 3^{2}=2,\quad 4^{2}=2,\quad 5^{2}=4,\quad 6^{2}=1,$
i.e., precisely the difference set in $\mathbb{Z}_{7}$ (i.e., the additive group of $\mathbb{F}_{7}$) in Example 1.9.
Thus, it follows from Theorem 1.10 that if $q\equiv 3\bmod 4$, $F$ is the Fourier transform of the additive group of $\mathbb{F}_{q}$, and $\Phi$ is the $(q-1)/2\times q$ submatrix of $F$ consisting of the rows indexed by $\operatorname{QR}(\mathbb{F}_{q})$, then $\Phi$ must be an ETF, which we call a Paley ETF. If $q\equiv 3\bmod 4$, we will denote by $\Phi_{q}$ this corresponding ETF.
We note that some authors (see, e.g., [BFMW13]) use the terminology “Paley ETF” to refer to $(d,2d)$-ETFs which are also constructed leveraging number theory about quadratic residues.
Example 2.3.
The smallest non-trival Paley ETF is $\Phi_{7}$. Namely, let $\zeta$ be a primitive $7$th root of unity. Then
$\Phi_{7}=\begin{pmatrix}1&\zeta&\zeta^{2}&\zeta^{3}&\zeta^{4}&\zeta^{5}&\zeta^{6}\\ 1&\zeta^{2}&\zeta^{4}&\zeta^{6}&\zeta&\zeta^{3}&\zeta^{5}\\ 1&\zeta^{4}&\zeta&\zeta^{5}&\zeta^{2}&\zeta^{6}&\zeta^{3}\end{pmatrix}$
is one ordering of the $(3,7)$-ETF which is the Paley ETF for $\mathbb{F}_{7}$. Note that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ and thus $\operatorname{NR}(\mathbb{F}_{7})=\{3,5,6\}$. Further, other than the 00th column, half of the columns have exponents that are $\operatorname{QR}$ and half have $\operatorname{NR}$. Finally, note that (cf. Example 1.12)
$\Phi_{7}=\left\{\,M^{k}\mathbbm{1}\,\,\middle|\,\,k\in\{0,\,\ldots,\,6\}\,\right\},\quad\textrm{where}\quad M=\begin{pmatrix}\zeta^{2^{0}}&0&0\\ 0&\zeta^{2^{1}}&0\\ 0&0&\zeta^{2^{2}}\end{pmatrix}\quad\mathbbm{1}=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}.$
The properties seen in Example 2.3 generalize. First, we define two special operators.
Definition 2.4.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $a$ be a generator of $\operatorname{QR}(\mathbb{F}_{p})$. The Paley modulation $M_{\{p\}}$ is the $(p-1)/2\times(p-1)/2$ diagonal matrix
$M_{\{p\}}=\begin{pmatrix}\zeta^{a^{0}}&0&\ldots&0\\ 0&\zeta^{a^{1}}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\zeta^{a^{(p-3)/2}}\end{pmatrix}.$
For any $d\in\mathbb{N}$, the $d\times d$ (cyclic) translation matrix $T_{d}$ is
$T_{d}=\begin{pmatrix}0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ 0&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&0\end{pmatrix}$
For a given $p$, the sizes of interest will be $M_{\{p\}}$ and $T_{(p-1)/2}$; thus, when clear from the context, we will often simply write $M$ and $T$.
Proposition 2.5.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $a$ be a generator of $\operatorname{QR}(\mathbb{F}_{p})$. Then the Paley ETF $\Phi_{p}$ may be constructed as
$\Phi_{p}=\left\{\,M_{\{p\}}^{k}\mathbbm{1}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\}\,\right\}.$
Proof.
This trivially follows from Example 1.12 and the fact that
$\operatorname{QR}(\mathbb{F}_{p})=\left\{\,a^{\ell}\,\,\middle|\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}.$
Note that, unlike in Example 2.3, the ordering of rows indexed using the construction above are not necessarily the ordering as a subsequence of $(1,\,\ldots,\,p-1)$. Proposition 2.5 also implies that other than $\mathbbm{1}$, $(p-1)/2$ vectors in $\Phi_{p}$ have entries equal to $\zeta$ raised to some permutation of the set of the $\operatorname{QR}$ and $(p-1)/2$ vectors entries which have $\operatorname{NR}$ as the exponents.
In order to use Theorem 1.14, we need to characterize the triple products of the Paley ETF vectors.
Lemma 2.6.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Further let $\zeta$ be the primitive $p$th root of unity and $a$ the generator of $\operatorname{QR}(\mathbb{F}_{p})$ used to generate $\Phi_{p}$. Finally, set
$\alpha=\sum_{\ell=0}^{(p-3)/2}\zeta^{a^{\ell}}.$
Then
$\operatorname*{TP}(j,k,\ell)=\left\{\begin{array}[]{lr}\alpha^{3};&j-k,k-\ell,\ell-j\in\operatorname{QR}\\[2.0pt] \left|\alpha\right|^{2}\alpha;&\textrm{ two of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \left|\alpha\right|^{2}\overline{\alpha};&\textrm{ one of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \overline{\alpha}^{3};&\ j-k,k-\ell,\ell-j\in\operatorname{NR}\end{array}\right.$
Proof.
First, we note that
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\sum_{\ell=0}^{(p-3)/2}\zeta^{ja^{\ell}}\zeta^{-ka^{\ell}}=\sum_{\ell=0}^{(p-3)/2}\zeta^{(j-k)a^{\ell}}=\left\{\begin{array}[]{lr}(p-1)/2&j=k\\ \alpha&j-k\in\operatorname{QR}\\ \overline{\alpha}&j-k\in\operatorname{NR}\end{array}\right.$
since $(j-k)a^{\ell}\in\operatorname{QR}$ if and only if $j-k\in\operatorname{QR}$ and also $\operatorname{NR}=-\operatorname{QR}$. So, for distinct $j,k,\ell$, the desired result immediately follows. ∎
We prove one more lemma before getting to the main result.
Lemma 2.7.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Let $\alpha$ be as in Lemma 2.6. Define $\Phi_{\operatorname{QR}}$ to be any $(p-1)/2\times(p-1)/2$ matrix consisting of the columns of $\Phi_{p}$ corresponding to the quadratic residues. Similarly define $\Phi_{\operatorname{NR}}$ for the non-quadratic residues. Then the row sums of $\Phi_{\operatorname{QR}}$ are $\alpha$, $\Phi_{\operatorname{QR}}$ is invertible, and the row sums of $\Phi_{\operatorname{QR}}^{-1}$ are $1/\alpha$. Similarly, the row sums of $\Phi_{\operatorname{NR}}$ are $1/\overline{\alpha}$, $\Phi_{\operatorname{NR}}$ is invertible, and the row sums of $\Phi_{\operatorname{NR}}^{-1}$ are $1/\overline{\alpha}$.
Proof.
The fact that $\Phi_{\operatorname{QR}}$ and $\Phi_{\operatorname{NR}}$ have the claimed row sums follows from the proof of Lemma 2.6. $\Phi_{\operatorname{QR}}$ and $\Phi_{\operatorname{NR}}$ are invertible due to a result by Chebotarëv about the invertibility of square submatrices of prime order Fourier transform matrices [SL96]. For the remaining claim, we note that for $n\times n$ $A=B^{-1}$ where $B$ has constant row sum $\beta$, for each $i\in\{1,\ldots,n\}$
11 11 $=\sum_{j=1}^{n}\delta_{i,j}=\sum_{j=1}^{n}\left(\sum_{k=1}^{n}A_{i,k}B_{k,j}\right)=\sum_{k=1}^{n}A_{i,k}\left(\sum_{j=1}^{n}B_{k,j}\right)=\sum_{k=1}^{n}A_{i,k}\beta=\beta\sum_{k=1}^{n}A_{i,k},$
as desired. ∎
Theorem 2.8.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$. Define $M=M_{\{p\}}$ and $T=T_{(p-1)/2}$. Then
$\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}$
$\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{(p-1)/2}\cong\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,m\in\operatorname{QR}(\mathbb{F}_{p}),\,\,b\in\mathbb{F}_{p}\,\right\}$
is the automorphism group of both the vectors and the lines of the Paley ETF $\Phi_{p}$. The group acts double-homogeneously but not double-transitively on the vectors and the lines.
A note of warning for those accustomed to thinking of an affine group as a translation group semidirect product a scaling group, in particular those with a background in wavelet theory: The group of matrices $\{T^{\ell}\}$ which translate the entries of the vectors act like dilations in the correspondence to $\operatorname{AGL}(\mathbb{F}_{p})$.
Proof.
First, we note that the order of $M$ is $p$ and the order of $T$ is $(p-1)/2$. This also means that raising $M$ to elements of $\mathbb{F}_{p}$ is well-defined when interpreted mod $p$. Further, for $a$ the given generator of the quadratic residues used to define $M$, $k\in\{0,\,\ldots,\,p-1\}$, and $\ell\in\{0,\,\ldots,\,(p-3)/2\}$, we have
$T^{\ell}M^{k}T^{-\ell}=M^{ka^{-\ell}}$
Thus,
$\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{(p-1)/2},$
with corresponding group operation
$\left(k,\ell\right)\left(\tilde{k},\tilde{\ell}\right)=\left(k+\tilde{k}a^{-\ell},\ell+\tilde{\ell}\right),$
which may be rewritten as
$\left(k,a^{\ell}\right)\left(\tilde{k},a^{\tilde{\ell}}\right)=\left(k+\tilde{k}a^{-\ell},a^{-\ell}a^{-\tilde{\ell}}\right),$
which is isomorphic to
$\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,a\in\operatorname{QR}(\mathbb{F}_{p}),\,\,b\in\mathbb{F}_{p}\,\right\}.$
By Proposition 3.1 in [Kan69], since $\{M^{k}T^{\ell}\}$ is isomorphic to the given subgroup of $\operatorname{AGL}(\mathbb{F}_{q})$, it suffices to show that it is the automorphism group and that it acts transitively as the automorphism group in order to show that it acts double-homogeneously but not double-transitively on either the vectors or the lines.
Recall that we denote the automorphism group of the vectors as $G$ and the automorphism group of the lines as $[G]$. Since we will implicitly show that the line automorphisms are $G\times U(1)$, we will treat elements of $[G]$ as being in $\operatorname{U}((p-1)/2)$ rather than $\operatorname{PU}((p-1)/2)$.
Note that Proposition 2.5 implies $G$ and $[G]$ are transitive since the vectors in $\Phi_{p}$ are the orbit of $\mathbbm{1}$ under $\{M^{k}\}$. We now show that $G$ contains $\{M^{k}T^{\ell}\}$. Trivially, the subgroup $\{T^{\ell}\}$ fixes $\mathbbm{1}$. Thus, for an arbitrary vector $M^{\kappa}\mathbbm{1}$ in $\Phi_{p}$, we have
$M^{k}T^{\ell}M^{\kappa}\mathbbm{1}=M^{k}M^{\kappa a^{-\ell}}T^{\ell}\mathbbm{1}=M^{k+\kappa a^{-\ell}}\mathbbm{1}\in\Phi_{p}$
From this it follows that $[G]\supseteq G\supseteq\{M^{k}T^{\ell}\}$.
Now we want to show that there is nothing else in $[G]$. The stabilizer of $\mathbbm{1}\mathbbm{1}^{\top}$ in $\{M^{k}T^{\ell}\}$ is just $\{T^{\ell}\}$. Since $[G]$ is transitive, by the orbit stabilizer theorem, the index of the stabilizer of $\mathbbm{1}\mathbbm{1}^{\top}$ in $[G]$, denoted $[G]_{\mathbbm{1}\mathbbm{1}^{\top}}$, is $p$. The index of $\{T^{\ell}\}$ in $\{M^{k}T^{\ell}\}$ is $p$. Thus, $[G]$ strictly contains $\{M^{k}T^{\ell}\}$ if and only if the stabilizer group of $\mathbbm{1}\mathbbm{1}^{\top}$ strictly contains $\{T^{\ell}\}$. Let $U\in[G]_{\mathbbm{1}\mathbbm{1}^{\top}}$. We seek to show that it is $T^{\ell}$ for some $\ell$ and thus $[G]=\{M^{k}T^{\ell}\}$, which would then imply that $G=\{M^{k}T^{\ell}\}$. Let $\sigma\in S_{0,1,\ldots,p-1}$ be the permutation of the $\varphi_{j}\varphi_{j}^{\ast}$ induced by $U$ (which necessarily fixes 00). By Theorem 1.14, we know that such a $\sigma$ must map triple products with $\operatorname*{TP}(0,k,\ell)$ to those $\operatorname*{TP}(0,\sigma(k),\sigma(\ell))$ of the same value. Plugging $j=0$ into Lemma 2.6, we have
$\operatorname*{TP}(0,k,\ell)=\left\{\begin{array}[]{lr}\alpha^{3}&k\in\operatorname{NR},\ell\in\operatorname{QR},k-\ell\in\operatorname{QR}\\[2.0pt] \left|\alpha\right|^{2}\alpha&(k,k-\ell\in\operatorname{NR},\,\ell\in\operatorname{QR})\textrm{ or }(k,\ell\in\operatorname{NR},\,k-\ell\in\operatorname{QR})\textrm{ or }(k,\ell,k-\ell\in\operatorname{QR})\\[2.0pt] \left|\alpha\right|^{2}\overline{\alpha}&(k,k-\ell\in\operatorname{QR},\,\ell\in\operatorname{NR})\textrm{ or }(k,\ell\in\operatorname{QR},\,k-\ell\in\operatorname{NR})\textrm{ or }(k,\ell,k-\ell\in\operatorname{NR})\\[2.0pt] \overline{\alpha}^{3}&\ k\in\operatorname{QR},\ell\in\operatorname{NR},k-\ell\in\operatorname{NR}\end{array}\right.$
For case $\alpha^{3}$, for any choice of $k\in\operatorname{NR}$, there is always at least one $\ell\in\operatorname{QR}$ such that $k-\ell\in\operatorname{QR}$ and vice versa. Thus, $\sigma$ must map $\operatorname{QR}$ to $\operatorname{QR}$ and $\operatorname{NR}$ to $\operatorname{NR}$. Now consider arbitrary $k\in\operatorname{QR}$ and $\ell=-k\in\operatorname{NR}$. The triple product is $\left|\alpha\right|^{2}\overline{\alpha}$ precisely when $k-\ell=2k\in\operatorname{QR}$, i.e., when $2\in\operatorname{QR}$. This happens when $p\equiv-1\bmod 8$. Otherwise, when $p\equiv 3\bmod 8$, the triple product is $\overline{\alpha}^{3}$. Either way, once the ordering of $\operatorname{NR}$ is fixed to be the ordering of negative the $\operatorname{QR}$, $\sigma$ induces the same permutation of the two sets (of order at most $(p-1)/2$). To that end, let $\Phi_{\operatorname{QR}}$ be the $(p-1)/2\times(p-1)/2$ circulant matrix that has the $\operatorname{QR}$ columns in order
$M^{1}\mathbbm{1},\,M^{(p-3)/2}\mathbbm{1},\,M^{(p-5)/2}\mathbbm{1},\,\ldots,\,M^{3}\mathbbm{1},\,M^{2}\mathbbm{1}.$
Define $\Phi_{\operatorname{NR}}=\overline{\Phi_{\operatorname{QR}}}$, the component-wise complex conjugation. Then $\Phi_{\operatorname{NR}}$ is also circulant and each column corresponds to the negative of the parameter of the column in $\Phi_{\operatorname{QR}}$. By Lemma 2.7, $\Phi_{\operatorname{QR}}$ and $\Phi_{\operatorname{NR}}$ are invertible and their inverses have constant row sums. Let $\Pi$ denote the $(p-1)/2\times(p-1)/2$ permutation matrix representing the permutation of the quadratic residues (respectively, the non residues). We would like to show that $U$ is this permutation (up to unimodular constant). On the vector level, in order for $U$ to induce the given permutations, there must exist diagonal unitary matrices $D_{\operatorname{QR}}$ and $D_{\operatorname{NR}}$ such that
$U=\Phi_{\operatorname{QR}}D_{\operatorname{QR}}\Pi\Phi_{\operatorname{QR}}^{-1}=\Phi_{\operatorname{NR}}D_{\operatorname{NR}}\Pi\Phi_{\operatorname{NR}}^{-1}$
Since $U\mathbbm{1}\mathbbm{1}^{\top}U^{\ast}=\mathbbm{1}\mathbbm{1}^{\top}$, $D_{\operatorname{QR}}$ and $D_{\operatorname{NR}}$ must be unimodular multiples of the identity and can be ignored. Thus, we have
$\left(\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}\right)\Pi=\Pi\left(\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}\right).$
As product of circulant matrices, $\Phi_{\operatorname{NR}}^{-1}\Phi_{\operatorname{QR}}$ is circulant and diagonalizable by the Fourier matrix for $\mathbb{Z}_{(p-1)/3}$. Since $\Pi$ commutes with it, it must also be diagonalizable by the Fourier matrix and hence circulant. The only circulant permutation matrices are of the form $T^{\ell}$ and the product commutes:
$U=\Phi_{\operatorname{QR}}T^{\ell}\Phi_{\operatorname{QR}}^{-1}=\Phi_{\operatorname{QR}}\Phi_{\operatorname{QR}}^{-1}T^{\ell}=T^{\ell},$
as desired. ∎
This theorem should not be incredibly surprising, given that Paley designs are the only $2$-homogeneous but not $2$-transitive Hadamard designs [Kan69]. As we will see in Theorem 3.6, Paley skew-conference matrices also yield an infinite class of ETFs which are $3$-homogeneous but not $3$-transitive. In [Kan69], it is also noted that if a set has a $2$-homogeneous automorphism group that is not $2$-transitive, represented as a subgroup of $\operatorname{AGL}(\mathbb{F}_{p})$, then the stabilizer of 00 (for us $\mathbbm{1}$ or $\mathbbm{1}\mathbbm{1}^{\top})$ permutes elements corresponding to the quadratic residues and elements corresponding the non-quadratic residues, which is what we observed.
3 $k$-Homogeneous ETFs
In this section, we characterize $k$-homogeneous ETFs. Throughout this section, we assume $d>1$. In particular, this means that the ETFs cannot trivially contain repeated elements.
The following result is generally known, but we include a proof here for completeness.
Lemma 3.1.
The line automorphism group of orthonormal bases and simplex ETFs with $n$ vectors is isomorphic to $S_{n}$. If an ETF is a regular simplex ETF or an orthonormal basis, then the vector automorphism group is also isomorphic to $S_{n}$.
Proof.
Given any two ordered orthonormal bases $U_{1}$ and $U_{2}$ of $\mathbb{C}^{n}$, $U_{2}U_{1}^{-1}$ is a unitary matrix that maps the basis $U_{1}$ to the basis $U_{2}$. Thus, the automorphism group of the vectors and in turn the lines is $S_{n}$. A $(d,n)$-ETF has $n=d+1$ vectors if and only if all of the triple products are negative [FJKM18]. Thus, by Theorem 1.14, any permutation of the lines is induced by some element of the automorphism group of the lines. Finally, if the ETF is a regular simplex ETF, then the inner products are all negative and thus any permutation of the vectors is possible (see, e.g., [CW16]). ∎
We also include the following converse.
Lemma 3.2.
The only ETFs with $3$-transitive line automorphism groups are orthonormal bases and simplices. The only ETFs with $2$-transitive vector automorphism groups are orthonormal bases and regular simplex ETFs. Finally, the only real ETFs with $2$-homogeneous vector automorphism groups with are orthonormal bases and regular simplex ETFs.
Proof.
The first result is Theorem II.6 in [Kin19]. The second comes from the fact that if the vector automorphism group is $2$-transitive, then any inner product of distinct vectors is mapped to any other inner product. The third comes from the fact that for real vectors the inner product is symmetric, meaning that $2$-homogeneity implies that all of the inner products between distinct vectors are the same. ∎
Note that this gives an alternate proof that the vector automorphism group of the prime order Paley ETFs $\Phi_{p}$ is not $2$-transitive. We note that the following immediately follows from the definition of homogeneity, but we record it here as we will use it frequently.
Lemma 3.3.
If a group acts $k$-homogeneously on a set of size $n$, then it also acts $(n-k)$-homogeneously.
Contrast this with transitivity, where every $k$-transitive group action is also $(k-1)$-transitive and where the larger $k$ is, the less likely it is that a group acts $k$-transitively. We now have a final lemma before stating and proving the main result of this section.
Lemma 3.4.
Assume that $\Phi$ is an ETF with $n>3$ vectors where the inner products of any two distinct vectors is imaginary. Then the line automorphism group of $\Phi$ is $3$-homogeneous. Conversely, if $n\notin\{d,d+1,2d\}$, $\Phi$ cannot have $3$-homogeneous automorphism group.
Proof.
Let $j,k,\ell$ be distinct indices. Then note that
$\overline{\operatorname*{TP}(j,k,\ell)}=\overline{\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle}$ $=\left\langle\varphi_{k},\varphi_{j}\right\rangle\left\langle\varphi_{\ell},\varphi_{k}\right\rangle\left\langle\varphi_{j},\varphi_{\ell}\right\rangle=\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{j}\right\rangle=\operatorname*{TP}(j,\ell,k).$
If the inner products are all imaginary, for any set $\{j,k,\ell\}$ of distinct indices there is some ordering of the triple products which is a positive real number times $i$. It now follows from Theorem 1.14 that the automorphism group of the lines is $3$-homogeneous.
Similarly, we note that for any ETF and any set $\{j,k,\ell\}$ of distinct indices, there are only two possible triple products regardless of order, $\operatorname*{TP}(j,k,\ell)$ and $\overline{\operatorname*{TP}(j,k,\ell)}$. Thus, if the triple products are not all complex conjugates of each other, there will be two sets of three distinct indices which cannot be mapped to each other. This means that the triple products of a $3$-homogeneous frame are either real or imaginary. Note that if the triple products are all positive then the vectors are switching equivalent to a set with all positive inner products and thus have an invertible Gram matrix. But the only non-redundant ETFs are orthonormal bases, a contradiction. Thus, the triple products are all $-\alpha$ or all ±$\pm i\alpha$ for some $\alpha\geq 0$. A necessary condition for an ETF to have purely imaginary inner products (see, e.g., [ET00, Str08]) is that $n=2d$. Simplices and orthonormal bases uniquely yield triple products which are all negative or 00, respectively. ∎
In [BP05, HS12], it was proven that redundant equiangular tight frames which are in a certain sense optimally robust to three erasures (called $3$-uniform) must be a simplex or have purely imaginary inner products, which suggests a connection to being $3$-homogeneous.
In [IM22], it was proven that any redundant spanning set which has a doubly transitive line automorphism group must be an ETF. However, the result holds for doubly homogeneous ETFs with minor changes to the proof.
Proposition 3.5.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ be a frame (i.e., spanning set) with $n>\max\{2,d\}$. If $\Phi$ has a doubly homogeneous line automorphism group, then $\Phi$ is an ETF.
Proof.
For any $U$ in the line automorphism group, there exists $\sigma\in S_{n}$ and $(\eta_{j})^{n}_{j=1}\in\operatorname{U}(1)^{n}$ such that
$U\varphi_{j}=\eta_{j}\varphi_{\sigma(j)},\quad j=1,2,\ldots,n.$
By [LW65, Theorem 2A], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is $(k-1)$-transitive. Thus, if $n\geq 4$, then the doubly homogeneous line automorphism group must act $1$-transitively, meaning that the vectors are all equal norm. If $n=3$, then the doubly homogeneous line automorphism group must act $(3-2)$-homogeneously, i.e., $1$-transitively, by Lemma 3.3, and again the vectors must be equal norm.
Let $j\neq k$ and $\ell\neq m$. Since the line automorphism group is doubly homogeneous, there exists a $U\in\operatorname{U}(d)$ with an associated permutation $\sigma$ yielding $\{\sigma(j),\sigma(k)\}=\{\ell,m\}$ as sets. Since $j,k,\ell,m$ were arbitrary, we have that there exist $0\leq\alpha\leq\beta$ such that for any $j,k$, $\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(\beta-\alpha)\delta_{j,k}+\alpha$.
A frame is tight if and only if $\Phi\Phi^{\ast}=AI$ if and only if the Gram matrix $\Gamma=\Phi^{\ast}\Phi$ is $A$ times an orthogonal projection. This is the formulation we will use. Let $j,k,\ell,m$ be as above. Then
$\Gamma_{j,k}$ $=\left\langle\varphi_{k},\varphi_{j}\right\rangle=\left\langle U\varphi_{k},U\varphi_{j}\right\rangle$
$=\left\langle\eta_{k}\varphi_{\sigma(k)},\eta_{j}\varphi_{\sigma(j)}\right\rangle$
$=\eta_{k}\overline{\eta_{j}}\left\langle\varphi_{\sigma(k)},\varphi_{\sigma(j)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\Gamma_{\sigma(j),\sigma(k)}$
=$=\left\{\begin{array}[]{c}\overline{\eta_{j}}\eta_{k}\Gamma_{m,\ell}\\[2.0pt] \overline{\eta_{j}}\eta_{k}\Gamma_{\ell,m}\end{array}\right.$
Similarly,
$\left(\Gamma^{2}\right)_{j,k}$ $=\sum_{i=1}^{n}\left\langle\varphi_{i},\varphi_{j}\right\rangle\left\langle\varphi_{k},\varphi_{i}\right\rangle$
$=\sum_{i=1}^{n}\left\langle U\varphi_{i},U\varphi_{j}\right\rangle\left\langle U\varphi_{k},U\varphi_{i}\right\rangle$
$=\sum_{i=1}^{n}\left\langle\eta_{i}\varphi_{\sigma(i)},\eta_{j}\varphi_{\sigma(j)}\right\rangle\left\langle\eta_{k}\varphi_{\sigma(k)},\eta_{i}\varphi_{\sigma(i)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\sum_{i=1}^{n}\left\langle\varphi_{\sigma(i)},\varphi_{\sigma(j)}\right\rangle\left\langle\varphi_{\sigma(k)},\varphi_{\sigma(i)}\right\rangle$
$=\overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{\sigma(j),\sigma(k)}$
=$=\left\{\begin{array}[]{c}\overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{m,\ell}\\[2.0pt] \overline{\eta_{j}}\eta_{k}\left(\Gamma^{2}\right)_{\ell,m}\end{array}\right.$
Thus, the off-diagonal of $\Gamma^{2}$ is a constant multiple of the off-diagonal of $\Gamma$. Also,
$\left(\Gamma^{2}\right)_{j,j}=\sum_{i=1}^{n}\left|\langle\varphi_{i},\varphi_{j}\rangle\right|^{2}=\beta^{2}+(n-1)\alpha^{2}=\frac{\beta^{2}+(n-1)\alpha^{2}}{\beta}\Gamma_{j,j}$
Hence, there exists $c_{1},c_{2}\in\mathbb{C}$ such that $\Gamma^{2}=c_{1}\Gamma+c_{2}I$. Since $n>d$, the Gram matrix $\Gamma$ cannot be invertible, meaning that $c_{2}=0$. (Otherwise $(\Gamma-c_{1}I)/c_{2}$ would be the inverse.) But this means that $\Gamma^{2}=c_{1}\Gamma$, meaning $\Gamma$ is a scalar multiple of an orthogonal projection and hence $\Phi$ is a tight frame. Putting everything together, we have that $\Phi$ is an ETF. ∎
In a sense, $2$-homogeneous ETFs are the only interesting $k$-homogeneous ETFs to study. $k$-homogeneous ETFs for $2<k\leq 5$ are either $2$-transitive (and thus one could use those tools to study them, like [IM22, IM24]) or in some sense either trivial or rare. We make this more explicit in what follows. The $(2,4)$-SIC is the (unique up to equivalence) ETF of $4$ vectors in $\mathbb{C}^{2}$.
Theorem 3.6.
The only ETFs with $3$-homogenous line automorphism groups are the $(2,4)$-SIC, the $(d,2d)$-ETFs with $d=q+1$ where $q\equiv 3\bmod 4$ is a prime power arising from Paley skew-conference matrices, orthonormal bases, and simplices, which are all $2$-transitive. The only ETFs with $4$-homogenous line automorphism groups are the $(2,4)$-SIC, the $(3,6)$ real ETF, orthonormal bases, and simplices. The only ETFs with $5$-homogenous line automorphism groups are the $(3,6)$ real ETF, the $(3,7)$- and $(4,7)$-ETFs (at least of Paley type), the $(4,8)$-ETF, orthonormal bases, and simplices.
Proof.
We consider $(d,n)$-ETFs, using [FM16] as a reference for which $(d,n)$ yield ETFs. By [LW65], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is $(k-1)$-transitive. Thus, all $3$-homogeneous ETFs (lines or vectors) with $n\geq 6$ are also $2$-transitive. By Lemma 3.4, the only ETFs which are not simplices or orthonormal bases but have $3$-homogeneous line automorphism group must have $n=2d$. It follows from [IM24] that the only $2$-transitive ETFs with $n=2d$ and purely imaginary inner products have $d=q+1$ with $q\equiv 3\bmod 4$ and are generated by Paley skew-conference matrices. The only ETF with $n<6$ which is not an orthonormal basis or simplex is the $(2,4)$ SIC, which is $1$-homogeneous as a SIC and thus $3=(4-1)$-homogeneous as well. Note that the $(2,4)$-SIC has $2$-transitive line automorphism group [Zhu15]; thus, all of the ETFs with $3$-homogeneous line automorphism group are also $2$-transitive.
Applying [LW65] again, we see that all $4$-homogeneous ETFs (lines or vectors) with $n\geq 8$ are also $3$-transitive, which means they are simplex ETFs or orthonormal bases (Lemma 3.2). Similarly, ETFs with $84\leq n<$ 8 which are orthonormal bases or simplex ETFs trivially have $4$-homogeneous line automorphism groups (Lemma 3.1). The remaining ETFs which have $n<8$ are the $(2,4)$-SICs, the $(3,6)$ real ETF (which is unique up to equivalence in $\mathbb{R}^{3}$ [BK06]), the $(3,7)$-ETF(s) and $(4,7)$-ETF(s). The $(2,4)$ SIC is trivially $4$-homogeneous since the identity maps the group of four lines to the group of four. The $(3,6)$ real ETF corresponds to the antipodal vertices of the regular icosahedron and thus has line automorphism group isomorphic to $\operatorname{PSL}(2,5)$, which acts $2$ transitively on the lines (see, e.g., Theorem 1.3 [IM24] which also gives uniqueness of the lines), in turn implying that the automorphism group is $2$-homogeneous and thus $4=(6-2)$-homogeneous. The remaining ETF parameters are the $(3,7)$- and $(4,7)$-ETFs. We have already shown that neither is $3$-homogeneous; thus, neither can have a $4=(7-3)$-homogeneous line automorphism group.
We now analyze ETFs for $5$-homogeneous line automorphism groups. As before, we only need to consider ETFs with $n<10=2(5)$. We have already shown that $(3,9)$ and $(6,9)$-ETFs cannot be $4$-homogeneous; thus, they are not $5=(9-4)$-homogeneous. Next we consider the $(4,8)$ ETF (which is $2$-transitive and unique up to equivalence [IM24]). In the standard construction, the inner products are all purely imaginary; thus, Lemma 3.4 implies that the ETF is $3$-homogeneous and in turn $5=(8-3)$-homogeneous. The $(3,7)$-Paley ETF and its Naimark complement are $2$-homogeneous, hence $5=(7-2)$-homogeneous. Finally, the $(3,6)$ real ETF is $1$-transitive, hence $1$-homogeneous and $5=(6-1)$-homogeneous. ∎
In the remaining cases where $k>5$, any $k$-homogeneous ETF with $n\geq 2k$ will be $(k-1)$-transitive and thus can only be a simplex ETF or orthonormal basis. For $n<2k$, the ETF will be $(n-k)$-homogeneous with $n-k<k$. Thus, inductively, the only possibilities are $1$-homogenous ETFs (i.e., the ETF lines are spun up as the orbit of a group with no additional symmetries), $2$-homogeneous ETFs, or any of the $3$, $4$, or $5$-homogeneous ETFs enumerated above.
Corollary 3.7.
The only ETFs with $3$-homogenous vector automorphism groups are orthonormal bases and the regular simplex ETFs. The only ETFs with $4$-homogenous vector automorphism groups are the $(2,4)$-SIC, orthonormal bases, and simplices.
Proof.
By following the cases in the proof of Theorem 3.6, we see that the only possible ETF which has a $3$-homogeneous vector automorphism group which is not doubly transitive is the $(2,4)$-SIC. Although any SIC is the orbit of a fiducial under a unitary system, the unitary system is a projective unitary representation and thus does not act transitively on the vectors. Hence, the vector automorphism group of the $(2,4)$-SIC is neither $1$-homogeneous nor $3=4-1$-homogeneous (Lemma 3.3). All other ETFs with ETFs with $3$-homogenous line automorphism groups must be doubly transitive, and hence Lemma 3.2 applies.
By Theorem 3.6, we just need to check the $(2,4)$-SIC and the $(3,6)$ real ETF for $4$-homogeneous automorphism vector groups. The $(2,4)$ SIC is trivially $4$-homogeneous. If the $(3,6)$ real ETF had a $4$-homogeneous vector automorphism group, then (Lemma 3.3) it would be $2=(6-4)$-homogeneous, which (Lemma 3.2 cannot happen since it is neither an orthonormal basis or a regular simple ETF. ∎
4 Dave Larson’s AMS Memoirs
Many infinite frames of note—like Gabor/time-frequency and wavelet—are generated as the (subsampling) of the orbit of a (projective) unitary representation and were a main focus of Dave Larson’s two AMS Memoirs [HL00, DL98]. A Gabor frame with time-frequency shifts over a lattice would be an example of an infinite frame with a transitive automorphism group of lines. We consider for a moment automorphism groups of such frames over infinite dimensional (separable) Hilbert spaces. Rephrasing Theorem 3.8 in [HL00]: If a tight frame $\Phi$ is the orbit of a unitary representation of a group $H$, then the Naimark complement is also the orbit of some unitary representation of the same group. Such a group $H$ is a subgroup of the automorphism groups of the lines and the vectors; so, this result is related to the fact that Theorem 1.14 implies that the automorphism groups of a finite tight frame and a Naimark complement are the same. Also, Corollary 3.10 of [HL00] yields that if a tight frame with frame bound $1$ is the orbit of a single bounded unitary operator, then it is unitarily equivalent to $\{e^{2\pi in\cdot}|_{E}\,:\,n\in\mathbb{Z}\}$ for $E\subset\mathbb{T}$, which is related to submatrices of Fourier matrices always yielding tight frames.
Assume that $\Phi=(\varphi_{j})_{j\in J}$ is a tight frame with frame bound $A$ for an infinite dimensional separable Hilbert space $\mathcal{H}$. Then $J$ is countably infinite. If $\Phi$ were doubly homogeneous, then $\left|\langle\varphi_{i},\varphi_{j}\rangle\right|=\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|$ for all $i\neq j$ and $k\neq\ell$. Call this number $\alpha$. By the definition of a tight frame (Definition 1.1), we see that
$\left\|\varphi_{k}\right\|^{2}=\frac{1}{A}\sum_{j\in J}\left|\langle\varphi_{k},\varphi_{j}\rangle\right|^{2}=\frac{1}{A}\left(\left\|\varphi_{k}\right\|^{2}+\sum_{j\neq k}\alpha^{2}\right).$
Since the left-hand-side is a finite number, the right-hand-side must converge, meaning that $\alpha=0$. Thus, the only infinite doubly homogeneous infinite frames are orthogonal bases.
An important concept introduced in [DL98] (with additional contemporary publications [DL98, HWW96, HWW97, FW96, DLS97, DLS98]) is wavelet sets. (Parseval) wavelet sets are measurable subsets of $\mathbb{R}^{d}$ for which the integer (lattice) translates satisfy one condition and the dyadic dilations satisfy another condition. When $X$ is a (Parseval) wavelet set, then the inverse Fourier transform of the indicator function of $X$ generates a (wavelet tight frame with frame bound $1$) a wavelet orthonormal basis. The simple geometric characterization of wavelet sets allowed the discovery of wavelets with surprising features. We may view the construction of Fourier ETFs as a kindred concept. By instead zeroing out rows rather than deleting them, we could restate Theorem 1.10 as meaning that modulating an indicator vector of a set yields a ETF for the span if and only if the set is a difference set in the group used to generate the Fourier matrix. As in Example 1.21, taking all of the shifts of any difference set (called the development) always yields a BIBD.
Finally, we note that one of the key tools developed in [DL98] was local (to a frame vector) commutants. This led to the analysis of the von Neumann algebra generated by the orbit of the left regular representation of a group $G$ being a tool to understand whether other representations of $G$ can yield frames [HL00]. A related idea is the use of Schurian association schemes to understand finite ETFs in [IM22, IM24]. It is possible that the similarities between these approaches could be leveraged to create further tools to characterize ETFs.
4.1 Problems
In honor of the “Problems” offered in Dave Larson’s Memoirs [HL00, DL98], we consider a handful of questions that this manuscript leads to.
When $q\equiv 3\bmod 4$ is a prime power $q=p^{s}$, one may still create a Paley ETF $\Phi_{q}$ by pulling out rows corresponding to the quadratic residues. As in Example 1.12, the frame vectors may be expressed as an orbit of $\mathbbm{1}$, i.e., $\varphi_{j}=M_{j}\mathbbm{1}$ for $M_{j}$ the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. However, when $s>1$, $\{M_{j}\}_{j}$ is a non-cyclic group (a representation of the non-cyclic additive group of $\mathbb{F}_{q}$). A certain group of translations will also be in the line and vector automorphism groups. In code from Joey Iverson, Harley Meade computed that the vector automorphism group of $\Phi_{3^{3}}$ is $(\mathbb{Z}_{3})^{3}\rtimes\mathbb{Z}_{3}$ while the line automorphism group is $\left((\mathbb{Z}_{3})^{3}\rtimes\mathbb{Z}_{13}\right)\rtimes\mathbb{Z}_{3}$. Calculations for larger non-cyclic Paley ETFs proved to be too unwieldy.
Question 4.1.
What are the vector and line automorphism groups of the Paley ETFs $\Phi_{q}$ with non-cyclic additive groups?
As referenced throughout this paper, various combinatorial designs arising from Paley-like constructions have reasonably high amounts of symmetry. It seems reasonable to assume that the non-cyclic Paley ETFs will be doubly homogeneous. We know that they are not doubly transitive [IM22, IM24]. By [Kan69], any set of size $n\geq 4$ which has a doubly homogeneous but not doubly transitive automorphism group must have $n=q$ for $q\equiv 3\bmod 4$ a prime power $q=p^{s}$ with the automorphism group isomorphic to a subgroup of
$\operatorname{A\Gamma L}(\mathbb{F}_{q})=\left\{\,f:\mathbb{F}_{q}\rightarrow\mathbb{F}_{q}\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,m\in\mathbb{F}_{q}^{\times},\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q}),\,\,b\in\mathbb{F}_{q}\,\right\}$
and isomorphically containing
$\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=x+b,\,\,b\in\mathbb{F}_{q}\,\right\}.$
Putting all of this together, we conjecture an answer to Question 4.1.
Conjecture 4.2.
Let $\Phi_{q}$ be a Paley ETF with $q=p^{s}$. Then $\Phi_{q}$ has a line automorphism group isomorphic to
$\left(\left(\mathbb{Z}_{p}\right)^{s}\rtimes\mathbb{Z}_{(q-1)/2}\right)\rtimes\mathbb{Z}_{s}\cong\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q}),\,\,b\in\mathbb{F}_{q}\,\right\}$
Due to a result from Chebotarëv [SL96], the Paley ETFs $\Phi_{p}$ with $p$ prime are full spark, i.e., have spark $(p-1)/2+1=(p+1)/2$. Their bender is just trivially the set of all subsets of size $(p+1)/2$, which is a $t$-design for all $21\leq t\leq(p+1)/$ 2. Calculations from Harley Meade show that the spark of $\Phi_{3^{3}}$ is $8$ and the bender is a $(27,8,28)$-BIBD. Again, calculations for larger non-cyclic Paley ETFs were not tractable.
Question 4.3.
What are the spark and bender of the Paley ETFs $\Phi_{q}$ with non-cyclic additive groups?
Dangerously powered by one infinite class and one example outside of this class, we propose the following conjecture.
Conjecture 4.4.
Let $\Phi_{q}$ be a Paley ETF with $q=p^{s}$. Then the spark of $\Phi_{q}$ is $((p+1)/2)^{s}$ and the bender is a $(p^{s},((p+1)/2)^{s},\lambda)$-BIBD.
We note that the computation of the Gram matrix of $\Phi_{p}$ for $p$ prime in the proof of Lemma 2.6 shows that the $\Phi_{p}$ are neither of two other categories of ETFs which are in some sense doubly-transitive-adjacent, i.e., being DRACKN or roux [CGSZ16, IM22]. This leads to the following question.
Question 4.5.
Are there any ETFs which are simultaneously doubly homogeneous and DRACKN or doubly homogeneous and roux but not doubly transitive?
It is a fact that any linear isometry of $\mathbb{C}^{d}$ with respect to $\ell^{p}$ with $p\neq 2$ is a generalized permutation matrix, i.e., a product of a diagonal unitary and a permutation matrix [LS94]. It is also a fact that any element $U$ of the automorphism group of a $(d,n)$-Fourier frame necessarily satisfies $\left\|U\varphi_{j}\right\|_{p}=\left\|\varphi_{j}\right\|_{p}$ for all $1\leq p\leq\infty$ with the $\varphi_{j}$ being a spanning set for $\mathbb{C}^{d}$. Being an isometry for $p\neq 2$ on a spanning set is not sufficient to prove that it extends to an isometry on the entire space. However, the computed automorphisms in this paper are all generalized automorphisms.
Question 4.6.
Are the elements of the automorphism group of a Fourier ETF (or, more generally, a Fourier frame) always generalized permutations?
Acknowledgments
The author is especially grateful to Harley Meade, who ran code to help formulate hypotheses and explicitly proved that a group related to $\{M^{k}T^{\ell}\}$ acts doubly homogeneously but not doubly transitively on the Naimark complement of a prime order Paley ETF (which was replaced with Proposition 3.1 in [Kan69] in the proof of Theorem 2.8). The author would also like to thank Joey Iverson and Dustin Mixon for fruitful discussions and also thank Joey for providing GAP code which allowed the exploration of various examples. Most of all, the author is grateful to Dave Larson for introducing her to frames twenty-something years ago.