### v2

Abstract
We consider geometric and combinatorial characterizations of equiangular tight frames (ETFs), with the former concerning homogeneity of the vector and line symmetry groups and the latter the matroid structure. We introduce the concept of the bender of a frame, which is the collection of short circuits, which in turn are the dependent subsets of frame vectors of minimum size. We also show that ETFs with $k$-homogeneous line symmetry groups have benders which are $k$-designs. Paley ETFs are a known class of ETFs constructed using number theory. We determine the line and vector symmetry groups of all Paley ETFs and show that they are $2$-homogeneous. We additionally characterize all $k$-homogeneous ETFs for $k\geq 3$. Finally, we revisit David Larson’s AMS Memoirs Frames, Bases, and Group Representations [HL00] coauthored with Deguang Han and Wandering Vectors for Unitary Systems and Orthogonal Wavelets [DL98] coauthored with Xingde Dai with a modern eye and focus on finite-dimensional Hilbert spaces. Keywords: equiangular tight frames, ETFs, $k$-homogeneous, $k$-transitive, Paley difference sets, $t$-designs, matroid circuits MSC2020: 42C15, 20B25, 05B10, 05B35
This is dedicated to David Larson, a great mentor and researcher.
1 Introduction
Frames are generalizations of orthonormal bases that have found wide applications in signal and image processing, compressed sensing, phase retrieval, and even characterizing emergent phenomenon in neural networks [CK13, KC07, FR13, PHD20]. Frames were originally introduced in [DS52] as a tool to loosen the definition of Fourier series. Frames over finite dimensional Hilbert spaces (see, e.g., [Wal18]), in particular so-called equiangular tight frames (ETFs) are intimately related to open problems in quantum information theory (like Zauner’s conjecture [Zau99, Zau11]) and combinatorial design theory (like the Gillespie conjecture [Gil18] [^1] It is possible that this conjecture is folklore. See, e.g., p. 263 of [GR01].). In this paper, we introduce the concept of the bender of a frame, which is the collection of all linearly dependent subsets of frame vectors of minimum size (called short circuits). We prove algebraic and combinatorial properties of an infinite class of ETFs (Paley ETFs). We also characterize all ETFs which have $k$-homogeneous symmetry groups for $k\geq 3$. Finally, we re-examine two of Dave Larson’s fundamental manuscripts from the turn of the millennium [HL00, DL98] in the context of frames in finite dimensions.
The author first learned about frames in 2002 in Dave Larson’s Research Experiences for Undergraduates (REU) on Frames and Wavelets, which ran for many years and inspired numerous young mathematicians. She has also used his book Frames for undergraduates [HKLW07]—cowritten with Deguang Han, Keri Kornelson, and Eric Weber, which came out of the experience running the REU—when working with students from high school to master’s level. Almost a quarter of a century later, the REU run by Dave Larson still has a large impact on her research program (and the research of many other alumni of the REU).
The basics of equiangular tight frames are covered in Section 1.1, followed by sections on their symmetries Section 1.2 and linear dependencies Section 1.3. In Definition 1.22 in Section 1.3, a new combinatorial concept is introduced: the bender, which consists of the short circuits of an ETF, which are the dependent subsets of minimum size. Section 2 contains results concerning Paley ETFs, with the main results being Theorems 2.10 and 2.19, in which the symmetry groups of all Paley ETFs are given. Section 3 concerns $k$-homogeneous ETFs, including Theorem 3.5 which characterizes $k$-homogeneous ETFs for $k\geq 3$. Finally, some key results in Dave Larson’s Memoirs [HL00, DL98] are contextualized within the framework of modern research on finite frames in Section 4. This section also includes a list of some questions and conjectures.
1.1 Equiangular Tight Frames
For any separable Hilbert space $\mathcal{H}$ and any orthonormal basis $\Phi=(\varphi_{j})_{j\in J}$, Parseval’s equality yields that for all $x\in\mathcal{H}$
$\left\|x\right\|^{2}=\sum_{j\in J}\left|\langle x,\varphi_{j}\rangle\right|^{2}.$
What might be surprising is that there are collections of vectors $\Phi$ which are not orthonormal bases which also satisfy that equality exactly (Parseval frames), in a scaled version (tight frames), or in a loosened version (frames). In infinite dimensions, even being a frame requires significant structure, and there is an immense amount of research about such vectors (see, e.g., [Chr16]).
Definition 1.1.
For a separable Hilbert space $\mathcal{H}$ and $\Phi=(\varphi_{j})_{j\in J}$ in $\mathcal{H}$, $\Phi$ is a tight frame if there exists $A>0$ such that for all $x\in\mathcal{H}$
$A\left\|x\right\|^{2}=\sum_{j\in J}\left|\langle x,\varphi_{j}\rangle\right|^{2}.$
AA is called the frame bound.
In finite dimensions, any spanning set is a frame (in general, not tight), and one often asks for additional structure, like being an equiangular tight frame. In this section only, we will use the notation $\mathbb{F}$ (without a subscript) to mean $\mathbb{R}$ or $\mathbb{C}$.
Definition 1.2.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ for $d\leq n$. We call $\Phi$ equiangular if there exist $a\leq b$ with $b>0$ such that
$\left|\langle\varphi_{j},\varphi_{k}\rangle\right|=(b-a)\delta_{j,k}+a.$
If additionally, $\Phi$ is a tight frame, then we call $\Phi$ an equiangular tight frame (ETF) or a $(d,n)$-ETF.
We will employ the common abuse of notation using $\Phi$ to both mean the frame itself and the short fat matrix with columns the frame vectors. With that notation, $\Phi$ is a tight frame precisely when $\Phi\Phi^{\ast}=AI$. Equiangular tight frames are optimally robust to noise and up to two erasures [GKK01, SH03, HP04], are important in quantum information theory [Zau99, Zau11], and provide optimal projective codes [JKM19]. As mathematical objects which interact with a variety of fields, there are a number of conjectures about their existence or the lack thereof: Zauner’s conjecture [Zau99, Zau11], the Gillespie conjecture [Gil18], the $(d,2d)$ conjecture [FI25, IJM24], and the Fickus-Jasper conjecture [FJ19].
For many choices of $d$, $n$, $\mathbb{F}$, there does not exist an ETF. (See, e.g., [FM16].) When $n=d$, the ETFs are simply equal-norm orthogonal bases. Up to equivalence (see below) when $n=d+1$, the vectors span lines which go through vertices of a regular simplex.
Definition 1.3.
If $\Phi$ is an ETF of $d+1$ vectors in $\mathbb{F}^{d}$, we call it a simplex ETF. If further, the inner product of any two distinct vectors in a simplex ETF is negative (necessarily the same negative number since they are equiangular), then we call it a regular simplex ETF.
In a certain sense, every simplex ETF is equivalent to a regular simplex ETF. Let $\operatorname{U}(d)$ denote the $d\times d$ orthogonal matrices when $\mathbb{F}=\mathbb{R}$ and $d\times d$ unitary matrices when $\mathbb{F}=\mathbb{C}$. Then there is always a $U\in\operatorname{U}(d)$ and a choice of scalars $\eta_{j}\in\operatorname{U}(1)$ such that
$\left(U\varphi_{j}\eta_{j}\right)_{j=1}^{d+1}$
have real entries and point to vertices of a regular $d$-simplex centered at the origin. In general, one may ask if two sequences of vectors are fundamentally “the same”, i.e., span the same set of lines up to a global change of basis. Fortunately, there is a relatively simple test.
Definition 1.4.
Assume that $\Phi=(\varphi_{j})_{j=1}^{n},\Psi=(\psi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$. We say that the vectors are switching equivalent (or projectively unitarily equivalent or isometrically isomorphic) if there exists a $U\in\operatorname{U}(d)$ and a choice of scalars $\eta_{j}\in\operatorname{U}(1)$ such that
$\psi_{j}=U\varphi_{j}\eta_{j},\quad j=1,2,\ldots,n.$
We note that some authors also allow permutation of the indices in their definition of switching equivalence. The test for switching equivalence has been rediscovered in different forms over many years (e.g., [Sei76, GP77, Bre90, AFF11, CW16]).
Proposition 1.5.
Let $\Phi=(\varphi_{j})_{j=1}^{n},\Psi=(\psi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ be ETFs. They are switching equivalent if and only if for $j\neq k\neq\ell\neq j$
$\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle=\left\langle\psi_{j},\psi_{k}\right\rangle\left\langle\psi_{k},\psi_{\ell}\right\rangle\left\langle\psi_{\ell},\psi_{j}\right\rangle.$
We will use the notation $\operatorname*{TP}(j,k,\ell)=\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle$, where $\Phi$ will be clear from context.
Every $(d,n)$-ETF over $\mathbb{F}$ has an (in general, infinitely many) associated (nd,n)(n-d,n)-ETF over $\mathbb{F}$ called a Naimark complement. Naimark complements of Parseval frames (not in general equiangular and allowing infinite dimensional Hilbert spaces) were first studied by Han and Larson [HL00], where they were called complementary frames. It was later realized that these results followed from Naimark’s dilation theorem concerning the lifting of positive operator valued measures [Neu43].
Proposition 1.6.
Let $\Phi$ be a $(d,n)$-ETF with $d<n$ over $\mathbb{F}$ with frame bound $A$. Then any $\Psi\in\mathbb{F}^{(n-d)\times n}$ that satisfies
$\Psi^{\ast}\Psi=AI-\Phi^{\ast}\Phi$
is an (nd,n)(n-d,n)-ETF over $\mathbb{F}$ with frame bound $A$ called a Naimark complement of $\Phi$.
Note that the inner products of distinct vectors of a Naimark complement are simply the negatives of the corresponding vectors in the first frame. Thus, by Proposition 1.5, characterizing switching equivalence of an ETF is equivalent to solving the problem for Naimark complements. Hence, when studying the switching equivalence of ETFs it suffices to study ETFs with $n\geq 2d$.
Given the historical connection of frames to Fourier bases [DS52], it is perhaps not surprising that Fourier-related concepts arise in constructions of various “nice” frames, like frames in $L^{2}(\mu)$ for $\mu$ a fractal measure (see, e.g., [DHSW11]) or Gabor frames for $L^{2}(\mathbb{R}^{d})$ (see, e.g., [Grö01]). Discrete Fourier transform matrices may also be used to define “nice” frames in finite dimensions.
Definition 1.7.
Let $\mathbb{Z}_{m}$ denote the integers mod $m$, and let $\zeta_{m}$ be a primitive $m$th root of unity. The Fourier matrix of $\mathbb{Z}_{m}$ is the $m\times m$ matrix with $(i,j)$-entry $\zeta^{ij}$, where indexing is $0,1,\ldots,m-1$. For $m_{1},\,m_{2},,\,\ldots,\,m_{k}\in\{2,\,3,\,\ldots\,\}$, the Fourier matrix of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$ is the Kronecker product of the Fourier matrices of the cyclic factors.
Note that this construction via Kronecker products yields a natural way to view the rows (and columns) of the Fourier matrix as being indexed lexicographically by the elements of $\mathbb{Z}_{m_{1}}\times\mathbb{Z}_{m_{2}}\times\cdots\times\mathbb{Z}_{m_{k}}$. By the fundamental theorem of finitely generated abelian groups, the character table of any finite abelian group is (up to possible permutations of the rows and columns) a Fourier matrix. Any $n\times n$ Fourier matrix is a unitary matrix scaled by $\sqrt{n}$; thus, any $d$-row all-column (i.e., $n$-column) submatrix $\Phi$ of an $n\times n$ Fourier matrix satisfies $\Phi\Phi^{\ast}=nI_{d}$. So, such submatrices always yield tight frames. We can completely characterize when such submatrices yield ETFs using combinatorial design theory (see, e.g., [CD07]).
Definition 1.8.
Let $G$ be a finite abelian group of size $v$ written additively. A $(v,k,\lambda)$-difference set is a subset $X\subseteq G$ of size $k$ such that the multiset of differences contains each non-identity element exactly $\lambda$ times.
Example 1.9.
Consider $\{1,2,4\}\subset\mathbb{Z}_{7}$. Then the difference table is
$\begin{array}[]{c||c|c|c}-&1&2&4\\ \hline\cr\hline\cr 1&0&1&3\\ \hline\cr 2&6&0&2\\ \hline\cr 4&4&5&0\end{array}$
That is, the multiset of differences contains each non-identity element of $\mathbb{Z}_{7}$ exactly once. Thus, $\{1,2,4\}$ is a $(7,3,1)$-difference set.
In any abelian group of size $v$, any subset of size $1$, v1v-1, or $v$ trivially forms a difference set. The difference set in Example 1.9 is the smallest nontrivial difference set. In particular, not all abelian groups have a nontrivial difference set. Examples of the construction of ETFs using difference sets appeared in [SH03, DF07, GR09], while the following theorem which completely characterizes the construction may be found in [XZG05].
Theorem 1.10.
Let $G$ be a finite abelian group with Fourier matrix $F$. Let $\Phi$ be an all-column submatrix of $F$, with rows indexed by $X$. Then $\Phi$ is an ETF if and only if $X$ is a difference set in $G$.
1.2 The Geometric Structure of ETFs
One way to further understand ETFs is by characterizing their geometric and combinatorial properties via their symmetries and linear dependencies. We introduce the former concept in this section and the latter in the next section.
Definition 1.11.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a frame for $\mathbb{F}^{d}$. We call $U\in\operatorname{U}(d)$ a vector automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}=\varphi_{\sigma(j)}$ for $j=1,\ldots n$. We call $U\in\operatorname{U}(d)$ a line automorphism of $\Phi$ if there exists a permutation $\sigma\in S_{n}$ such that $U\varphi_{j}\varphi_{j}^{\ast}U^{\ast}=\varphi_{\sigma(j)}\varphi_{\sigma(j)}^{\ast}$ for $j=1,\ldots n$. The subgroup of $S_{n}$ induced from the vector automorphisms is called the vector symmetry group $\operatorname*{Sym}_{v}(\Phi)$ of $\Phi$, and the subgroup of $S_{n}$ induced from the line automorphisms is called the line symmetry group $\operatorname*{Sym}_{\ell}(\Phi)$ of $\Phi$.
Example 1.12.
Let $G$ be a finite abelian group with $n\times n$ Fourier matrix $F$. Let $\Phi$ be a $d$-row, all-column submatrix of $F$, which is necessarily a tight frame. Further, define $M^{\{j\}}$ for each frame index $j$ as the $d\times d$ diagonal matrix with diagonal entries the entries of $\varphi_{j}$. Then $\{M^{\{j\}}\}_{j}$ is a subgroup of $\operatorname{U}(d)$ which is a representation of $G$. Also, $F$ is the orbit of the all-ones vector $\mathbbm{1}$ under $\{M^{\{j\}}\}_{j}$, where $M^{\{j\}}\mathbbm{1}=\varphi_{j}$. This in turn implies that $\operatorname*{Sym}_{v}(\Phi)\leq\operatorname*{Sym}_{\ell}(\Phi)$ have a subgroup isomorphic to $G$.
We will often ask when the automorphism groups have extra structure.
Definition 1.13.
An action of a group $G$ on a non-empty set $X=\left\{\,x_{i}\,\,\middle|\,\,i\in\{1,\ldots,n\}\,\right\}$ is $k$-transitive if for each ordered $k$-tuple with no repeated elements $(x_{i_{1}},\ldots,x_{i_{k}})$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as ordered tuples)
$(g\cdot x_{i_{1}},\,g\cdot x_{i_{2}},\,\ldots,\,g\cdot x_{i_{k}})=(x_{\sigma(i_{1})},\,x_{\sigma(i_{2})},\,\ldots,\,x_{\sigma(i_{k})}).$
The action is $k$-homogeneous if for each set of $k$ elements of $X$ with no repeated elements $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ and each permutation $\sigma\in S_{n}$, there exists a $g\in G$ such that (as sets)
$\{g\cdot x_{i_{1}},\,g\cdot x_{i_{2}},\,\ldots,\,g\cdot x_{i_{k}}\}=\{x_{\sigma(i_{1})},\,x_{\sigma(i_{2})},\,\ldots,\,x_{\sigma(i_{k})}\}.$
A frame is $k$-transitive (respectively, $k$-homogeneous) if its line automorphism group is $k$-transitive (respectively, $k$-homogeneous).
Clearly, $k$-transitivity implies both (k1)(k-1)-transitivity and $k$-homogeneity. However, the relationship between $k$-homogeneity and $\ell$-homogeneity for $k\neq\ell$ is subtler. Consider a set of size $k$ acted on by the trivial group. Then the group action is $k$-homogeneous but not $j$-homogeneous for any $1\leq j<k$. On the other hand, when $k$ is less than or equal to half the size of the set acted on, $k$-homogeneity implies both (k1)(k-1)-transitivity and (k1)(k-1)-homogeneity [DM96, LW65] and in many cases also $k$-transitivity [Kan72]. In Section 2, we will prove that an infinite class at ETFs have symmetry groups which are $2$-homogeneous but not $2$-transitive.
We make a note of a result from [CW18] (cf. Proposition 1.5) that we will leverage to prove a few results and which was also the basis of the code [Ive] provided by Joey Iverson that was used to test some ideas.
Theorem 1.14.
[CW18] Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. Then $\sigma\in\operatorname*{Sym}_{v}(\Phi)$ if and only if for all $j\neq k$
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\left\langle\varphi_{\sigma(j)},\varphi_{\sigma(k)}\right\rangle.$
Similarly, $\sigma\in\operatorname*{Sym}_{\ell}(\Phi)$ if and only if for all $j\neq k\neq l\neq j$
$\operatorname*{TP}(j,k,\ell)=\operatorname*{TP}(\sigma(j),\sigma(k),\sigma(\ell)).$
Since this result relies solely on the inner products, one can see that the vector symmetry groups (respectively, the line symmetry groups) of an ETF and a Naimark complement of an ETF are the same.
1.3 The Combinatorial Structure of ETFs
Understanding the linear dependencies of vectors in ETFs rose to prominence in the search for so-called deterministic restricted isometry property (RIP) measurements in compressed sensing (see, e.g., [BFMW13]). However, it was realized that many ETFs have in some sense worst possible linear dependence properties [FMT12].
Definition 1.15.
Let $\Phi=(\varphi_{j})_{j=1}^{n}\in\mathbb{F}^{d\times n}$ for $n>d$. The spark (or girth) of $\Phi$, $\operatorname{spark}(\Phi)$, is the size of the smallest subset of linearly dependent vectors in $\Phi$.
“Spark” is the term from compressed sensing literature [DE03], while “girth” is a more classical term from matroid theory (see, e.g., [Oxl11]). RIP concerns subsets up to a certain size of columns of a matrix having condition number close to $1$; so, small subsets of columns being linearly dependent means that the matrix is far from being RIP.
Example 1.16.
Consider an orthonormal basis $(u_{j})_{j=1}^{d}$ for $\mathbb{F}^{d}$. Then
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&\sum_{j=1}^{d}u_{j}\end{pmatrix}=d+1$,
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&u_{1}\end{pmatrix}=2$, and
  • $\operatorname{spark}\begin{pmatrix}u_{1}&\ldots&u_{d}&0\end{pmatrix}=1$,
where the singleton set consisting of the zero vector is counted as a linearly dependent set of size $1$, while the rank of each of the vector collections above is $d$.
Trivially, the largest that the spark of $n>d$ vectors in $\mathbb{F}^{d}$ may be is $d+1$, which happens when any subset of $d$ vectors is linearly independent. When this happens, we say the vectors are full spark. Full spark frames (respectively, full spark tight frames) of $n$ vectors in $\mathbb{F}^{d}$ form an open dense set in the space of frames (respectively, tight frames) of $n$ vectors in $\mathbb{F}^{d}$ [Cah09, ACM12, CMS17]. However, the same cannot be said when restricting to ETFs. Applying a now classical result from the compressed sensing literature [DE03] to ETFs, we have the following result.
Proposition 1.17.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a $(d,n)$-ETF with $n>d$. Then
$\operatorname{spark}(\Phi)\geq 1+\sqrt{\frac{d(n-1)}{n-d}}.$
The infinite class of ETFs constructed in [FMT12] all saturate this worst-case bound for spark. Inspired by this, the author of this paper and coauthors characterized all ETFs which saturate this bound [FJKM18].
Theorem 1.18.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be a $(d,n)$-ETF with $n>d$. Then
$\operatorname{spark}(\Phi)=1+\sqrt{\frac{d(n-1)}{n-d}}$
if and only if $\Phi$ contains a subset of vectors which form a simplex ETF for their span.
This led to the following definition [FJKM18].
Definition 1.19.
Let $\Phi=(\varphi_{j})_{j=1}^{n}$ be an ETF. The binder of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which form simplex ETFs for their span.
If the binder is non-empty, then each of the elements of the binder (which are subsets of $\Phi$) have the same size. In certain cases (e.g., Gabor-Steiner frames [BK19]), the binder forms a type of combinatorial design call a BIBD. For more on combinatorial designs see, e.g., [CD07].
Definition 1.20.
A $t$-$(v,k,\lambda)$-design or block design is an ordered pair $(V,\mathcal{B})$, where $V$ is a set of size $v$ and $\mathcal{B}$ is a collection of subsets of $V$ called blocks such that each block $\beta\in\mathcal{B}$ has the same size $k$ and any subset of $t$ distinct elements in $V$ is contained in precisely $\lambda>0$ blocks. When $t=2$, the design is also called a balanced incomplete block design (BIBD).
Note that $t$-designs are also $s$-designs for any $1\leq s\leq t$.
Example 1.21.
Consider the following subsets of $\mathbb{Z}_{7}$ which result from shifting the difference set in Example 1.9 by each element of the group:
$\left\{\{1,2,4\},\,\{2,3,5\},\,\{3,4,6\},\,\{0,4,5\},\,\{1,5,6\},\,\{0,2,6\},\,\{0,1,3\}\right\}.$
Note that $\{1,2\}$ is only in the first block, $\{1,3\}$ only in the seventh, $\{1,4\}$ only in the first, and so on. This is a 22-$(7,3,1)$-design a.k.a. a $(7,3,1)$-BIBD.
In practice we find that many ETFs have an empty binder. This does not mean that the ETF is full spark, only that it does not have worst case spark. So, one may ask to study the properties of the smallest dependent sets. This leads us to introduce a new definition.
Definition 1.22.
Let $\Phi$ be an ETF with spark $s$. The bender222“Bender” is in honor of the smack-talking robot on Futurama, a show which is filled with math jokes. of $\Phi$ is the set of subsets of $\Phi$ (or their corresponding indices) which are size $s$ and linearly dependent. The elements of the bender are called short circuits.
In matroid theory, minimally dependent subsets of vectors, i.e., those for which any proper subset is independent, are called circuits. The size of the smallest circuit(s) is the girth, i.e., the spark. Hence the name “short circuit.” [^3] Thanks to Matt Fickus, who suggested this name.
If $\Phi$ has a nontrivial binder, then the bender is the binder and the short circuits are all simplex ETFs for their span. On the other extreme, if $\Phi$ is full spark $d+1$, then the bender is the set of all subsets of size $d+1$ of the vectors.
It was noted in [FL25] that the binder of an ETF with a $2$-transitive line automorphism group is a BIBD. However, folklore in combinatorial design theory (see, e.g., [CD07, Remark 4.29]) means a much more general result holds.
Proposition 1.23.
Let $\Phi$ be an ETF with $k$-homogeneous line symmetry group and spark $\geq k$. Then the bender forms a $k$-design.
Proof.
Note that neither scaling vectors by unimodular scalars nor performing a unitary change of basis affects linear dependency of a set. So, we will consider the short circuits to consist of the sets of the lines spanned by each of the vectors, and, thus, acting on a short circuit with a line automorphism results in a short circuit. For any two different sets $X_{1}$ and $X_{2}$ of $k$ frame vectors, there is a least one element of the line symmetry group mapping $X_{1}$ to $X_{2}$, meaning that the size of the set of short circuits containing $X_{1}$ is the same as the size of the set of short circuits containing $X_{2}$. Thus, the bender is a $k$-design. ∎
2 Paley ETFs
We will apply the tools from Sections 1.2 and 1.3 to combinatorially and geometrically characterize one infinite class of ETFs, the Paley (difference set) ETFs. These ETFs are generated using the structure of finite fields. Thus, we begin in Section 2.1 by reviewing some basic facts about finite fields. For each prime power $q\equiv 3\bmod 4$, there is a Paley ETF $\Phi_{q}$. We will prove in Sections 2.22.4 that $\operatorname*{Sym}_{v}(\Phi_{q})=\operatorname*{Sym}_{\ell}(\Phi_{q})$ act $2$-homogeneously but not $2$-transitively on the vectors and the lines, respectively. The Paley ETFs thus have a bender which is a BIBD. In Section 2.2, we prove the result (Theorem 2.10) when $q$ is prime as the proof requires less machinery and gives a flavor of the arguments for the more general case. This result also explicitly gives the unitary matrices which are vector automorphisms. Then in Section 2.3, we provide an alternate proof of Theorem 2.10 that filters through Theorem 1.14. This proof does not yield the explicit vector automorphisms but yields techniques which may be useful for other classes of ETFs. A quick example of the technique being applied to another infinite class of ETFs (Gabor-Steiner) is also given. Finally, in Section 2.4, we prove the result (Theorem 2.19) for all prime powers $q\equiv 3\bmod 4$.
2.1 Finite Fields
[…]
$\operatorname{QR}(\mathbb{F}_{q})$ $=\left\{\,x^{2}\,\,\middle|\,\,x\in\mathbb{F}_{q}^{\times}\,\right\}$
$\operatorname{NR}(\mathbb{F}_{q})$ $\QR.\displaystyle=\mathbb{F}_{q}^{\times}\backslash\operatorname${QR}.
Thus, $\mathbb{F}_{q}=\operatorname{QR}(\mathbb{F}_{q})\cup\operatorname{NR}(\mathbb{F}_{q})\cup\{0\}$. When the underlying field is clear from context, we will often simply write $\operatorname{QR}$ and $\operatorname{NR}$. Note that $\operatorname{QR}$ is thus a subgroup of the multiplicative cyclic group $\mathbb{F}_{q}^{\times}$ of index $2$, meaning it has size (q1)/2(q-1)/2 and is also cyclic and generated by a square of a primitive element.
Example 2.1.
We have that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ since in $\mathbb{F}_{7}$
$1^{1}=1,\quad 2^{2}=4,\quad 3^{2}=2,\quad 4^{2}=2,\quad 5^{2}=4,\quad 6^{2}=1,$
i.e., precisely the difference set in $\mathbb{Z}_{7}$ (i.e., the additive group of $\mathbb{F}_{7}$) in Example 1.9. Note further that $3$ and $5$ are primitive elements, i.e., $3>=<5>=\mathbb{F}_{7}^{\times}$ and that $2=3^{2}$ and $4=5^{2}$ generate $\operatorname{QR}(\mathbb{F}_{7})$.
For fields of prime order $p$, the structure is relatively simple. The additive group $(\mathbb{F}_{p},+)$ and the multiplicative group $(\mathbb{F}_{p}^{\times},\cdot)$ are both cyclic. For $q=p^{s}$ with $s>1$, the additive group $(\mathbb{F}_{q},+)$ is no longer cyclic. Further, we can think of $\mathbb{F}_{q}$ as both a field extension of $\mathbb{F}_{p}$ and a vector space over $\mathbb{F}_{p}$. We will need to leverage both types of structure in our analysis.
First, we represent $\mathbb{F}_{q}$ as a field extension. We pick a monic irreducible (and thus separable since $\mathbb{F}_{p}$ is a finite field) polynomial $\pi\in\mathbb{F}_{p}[X]$ of degree $s$. Then $\mathbb{F}_{q}$ is the Galois extension $\mathbb{F}_{q}\cong\mathbb{F}_{p}(\alpha)\cong\mathbb{F}_{p}[X]/(\pi(X))$, where $\pi(\alpha)=0$. Since $\mathbb{F}_{q}$ is also congruent to the splitting field of $X^{q}-X$, we can (and will) choose $\pi$ so that $\alpha$ is a primitive element and $\alpha^{2}$ generates $\operatorname{QR}$.
We may view the vector space structure of $\mathbb{F}_{q}$ as coordinate vectors in $\mathbb{F}_{p}^{s}$ via the primitive element basis $\alpha^{0},\,\alpha^{1},\,\ldots\alpha^{s-1}$. The Frobenius automorphism $x\mapsto x^{p}$ generates the Galois group of automorphisms. A basis which is the orbit of an element under the Galois group, i.e., $\left\{\,x^{p^{i}}\,\,\middle|\,\,i\in\{0,1,\ldots,s-1\}\,\right\}$ for particular $x\in\mathbb{F}_{q}$ is called a normal basis.
Example 2.2.
Consider $\pi(x)=x^{3}+2x+1\in\mathbb{F}_{3}[X]$. Then $\pi$ is irreducible over $\mathbb{F}_{3}$ since $0^{3}+2(0)+1,1^{3}+2(1)+1,2^{3}+2(2)+1\neq 0$. Let $\alpha$ denote a root of $\pi$. Then
$\begin{array}[]{ccccc}\alpha^{0}=1,&\alpha^{1}=\alpha,&\alpha^{2}=\alpha^{2},&\alpha^{3}=2+\alpha,&\alpha^{4}=2\alpha+\alpha^{2},\\ \alpha^{5}=2+\alpha+2\alpha^{2},&\alpha^{6}=1+\alpha+\alpha^{2},&\alpha^{7}=2+2\alpha+\alpha^{2},&\alpha^{8}=2+2\alpha^{2},&\alpha^{9}=1+\alpha,\\ \alpha^{10}=\alpha+\alpha^{2},&\alpha^{11}=2+\alpha+\alpha^{2},&\alpha^{12}=2+\alpha^{2},&\alpha^{13}=2,&\alpha^{14}=2\alpha,\,\ldots\end{array}$
where $\alpha^{13+k}=2\alpha^{k}$ for $k\in\{0,\ldots,13\}$. That is, $\alpha$ is a primitive element, $\{\alpha^{0},\alpha^{1},\alpha^{2}\}$ is a primitive element basis for $\mathbb{F}_{27}\cong\mathbb{F}_{3}(\alpha)$, and $\operatorname{QR}(\mathbb{F}_{27})=<\alpha^{2}>$. We also note that the orbit of $\alpha^{2}$ under the Galois group forms a normal basis. Namely,
$\alpha^{2},\quad\left(\alpha^{2}\right)^{3}=\alpha^{6}=1+\alpha+\alpha^{2},\quad\left(\alpha^{2}\right)^{9}=\alpha^{18}=2\alpha^{5}=1+2\alpha+\alpha^{2}$
form a basis for $\mathbb{F}_{27}$. One way to see this is by inspecting the coordinate vectors with respect to the primitive element basis:
$\begin{pmatrix}0&0&1\end{pmatrix}^{\top},\quad\begin{pmatrix}1&1&1\end{pmatrix}^{\top},\quad\begin{pmatrix}1&2&1\end{pmatrix}^{\top}$.
We denote the vector space of $s\times s$ matrices with entries in a field $\mathbb{F}$ as $\mathbb{F}^{s\times s}$. Let $C\in\mathbb{F}_{p}^{s\times s}$ be the companion matrix of $\pi$. Then the group generated by $C$ is isomorphic to $\mathbb{F}_{q}^{\times}$ via the map $C\mapsto\alpha$.
Example 2.3.
Consider the construction of $\mathbb{F}_{27}$ from Example 2.2. The companion matrix of $\pi$ in $\mathbb{F}_{3}^{3\times 3}$ is
$C=\begin{pmatrix}0&0&2\\ 1&0&1\\ 0&1&0\end{pmatrix}$.
Let $e_{1}$ be the first standard basis vector (i.e., $e_{1}$ represents the multiplicative identity in $\mathbb{F}_{27}$). Then, $Ce_{1}$ represents $\alpha$ using coordinates from the primitive basis $\{1,\alpha,\alpha^{2}\}$ and further
$\mathbb{F}_{27}^{\times}=\left\{\,C^{k}e_{1}\,\,\middle|\,\,k\in\{0,\ldots,26\}\,\right\}\quad\textrm{and}\quad\operatorname{QR}(\mathbb{F}_{27})=\left\{\,C^{2k}e_{1}\,\,\middle|\,\,k\in\{0,\ldots,12\}\,\right\}.$
Also, $\left\{\,C^{2k}e_{1}\,\,\middle|\,\,k\in\{1,3,9\}\,\right\}$ forms a normal basis for $\mathbb{F}_{27}$.
When $q\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$; that is, the quadratic non-residues are precisely the negations of the quadratic residues. This fact implies the following (see, e.g. [CD07]).
Theorem 2.4.
When $q\equiv 3\bmod 4$, $\operatorname{QR}(\mathbb{F}_{q})$ forms a (q,(q1)/2,(q3)/4)(q,(q-1)/2,(q-3)/4)-difference set in the additive group of $\mathbb{F}_{q}$.
Thus, it follows from Theorem 1.10 that if $q\equiv 3\bmod 4$, $F$ is the Fourier transform of the additive group of $\mathbb{F}_{q}$, and $\Phi$ is the $(q-1)/2\times q$ submatrix of $F$ consisting of the rows indexed by $\operatorname{QR}(\mathbb{F}_{q})$, then $\Phi$ must be an ETF, which we call a Paley ETF. If $q\equiv 3\bmod 4$, we will denote by $\Phi_{q}$ this corresponding ETF. We will fix an ordering of the rows of $\Phi_{q}$ as $\alpha^{0},\alpha^{2},\ldots,\alpha^{q-3}$ for a primitive element $\alpha$. This will allow the explicit definition of matrices in $\operatorname{U}((q-1)/2)$ which are the vector automorphisms. The ordering of the columns is less important than the parameterization, as we just need to characterize the induced permutations of $\mathbb{F}_{q}$.
We note that some authors (see, e.g., [BFMW13]) use the terminology “Paley ETF” to refer to $(d,2d)$-ETFs which are also constructed leveraging number theory about quadratic residues.
Example 2.5.
The smallest nontrivial Paley ETF is $\Phi_{7}$. Namely, let $\zeta$ be a primitive $7$th root of unity. Then
$\Phi_{7}=\begin{pmatrix}1&\zeta&\zeta^{2}&\zeta^{3}&\zeta^{4}&\zeta^{5}&\zeta^{6}\\ 1&\zeta^{2}&\zeta^{4}&\zeta^{6}&\zeta&\zeta^{3}&\zeta^{5}\\ 1&\zeta^{4}&\zeta&\zeta^{5}&\zeta^{2}&\zeta^{6}&\zeta^{3}\end{pmatrix}$
is one ordering of the $(3,7)$-ETF which is the Paley ETF for $\mathbb{F}_{7}$. Note that $\operatorname{QR}(\mathbb{F}_{7})=\{1,2,4\}$ and thus $\operatorname{NR}(\mathbb{F}_{7})=\{3,5,6\}$. Further, other than the 0th column, half of the columns have exponents that are $\operatorname{QR}$ and half have $\operatorname{NR}$. Finally, note that (cf. Example 1.12)
$\Phi_{7}=\left\{\,M^{k}\mathbbm{1}\,\,\middle|\,\,k\in\{0,\,\ldots,\,6\}\,\right\},\quad\textrm{where}\quad M=\begin{pmatrix}\zeta^{2^{0}}&0&0\\ 0&\zeta^{2^{1}}&0\\ 0&0&\zeta^{2^{2}}\end{pmatrix}\quad\mathbbm{1}=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$.
The ordering of the rows in this example corresponds to the choice of primitive element $3$, i.e.,
$\left(3^{2(0)},\,3^{2(1)},\,3^{2(2)}\right)=\left(2^{0},\,2^{1},\,2^{2}\right).$
Consider the following bilinear form on $\mathbb{F}_{p}^{s}$:
$\left\langle\begin{pmatrix}a_{0}\\ a_{1}\\ \vdots\\ a_{s-1}\end{pmatrix},\begin{pmatrix}b_{0}\\ b_{1}\\ \vdots\\ b_{s-1}\end{pmatrix}\right\rangle_{p}=\sum_{i=0}^{s-1}a_{i}b_{i}.$
Let $p$ be prime and $s\geq 0$. Further let $\zeta$ be a primitive $p$th root of unity. Up to possible reorderings of the rows and columns, the Fourier matrix of $(\mathbb{F}_{p}^{s},+)$ has entries $(\zeta^{\left\langle x,y\right\rangle_{p}})_{x,y\in\mathbb{F}_{p}^{s}}$. This follows immediately from the Kronecker product construction in Definition 1.7.
Example 2.6.
Continuing Examples 2.2 and 2.3, let $\zeta$ be a primitive third root of unity so that the Paley ETF $\Phi_{27}$ has entries
$\left\{\,\zeta^{\left\langle x,y\right\rangle_{p}}\,\,\middle|\,\,x\in\operatorname{QR},\,y\in\mathbb{F}_{27}\,\right\}.$
Since $Ce_{1}$ is a primitive element and $C^{2}e_{1}$ generates the quadratic residues, we may parameterize each column except the one indexed by $y=0$ (which has all entries equal to 11) as
$\left\{\,\zeta^{\left\langle x,y\right\rangle_{p}}\,\,\middle|\,\,x\in\operatorname{QR},\,y\in\mathbb{F}_{27}^{\times}\,\right\}=\left\{\,\zeta^{\left\langle C^{2k}e_{1},C^{\ell}e_{1}\right\rangle_{p}}\,\,\middle|\,\,k\in\{0,\ldots,12\},\,\ell\in\{0,\ldots,25\}\,\right\}.$
2.2 Paley ETFs over Fields of Prime Order
In this section, we will explicitly give the vector automorphisms of $\Phi_{p}$ for $p\equiv 3\bmod 4$ prime and show that the vector and line symmetry groups act $2$-homogeneously but not $2$-transitively. The line automorphism group is generated by the following operators.
Definition 2.7.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Further let $\alpha$ be a primitive element of $\mathbb{F}_{p}$. The (cyclic) Paley modulation $M_{\{p\}}$ is the $(p-1)/2\times(p-1)/2$ diagonal matrix
$M_{\{p\}}=\begin{pmatrix}\zeta^{\alpha^{0}}&0&\ldots&0\\ 0&\zeta^{\alpha^{2}}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\zeta^{\alpha^{p-3}}\end{pmatrix}$.
For any $d\in\mathbb{N}$, the $d\times d$ (cyclic) translation matrix $T_{d}$ is
$T_{d}=\begin{pmatrix}0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ 0&1&\ldots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\ldots&1&0\end{pmatrix}$.
For a given $p$, the sizes of interest will be $M_{\{p\}}$ and $T_{(p-1)/2}$; thus, when clear from the context, we will often simply write $M$ and $T$. As in Example 2.5, we may characterize Paley ETFs over prime-order fields as the orbit under the cyclic group of Paley modulations.
Proposition 2.8.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and $\zeta$ be a primitive $p$th root of unity. Then the Paley ETF $\Phi_{p}$ may be constructed as
$\Phi_{p}=\left\{\,M_{\{p\}}^{k}\mathbbm{1}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\}\,\right\}.$
Proof.
Let $\alpha$ be the primitive element of $\mathbb{F}_{p}$ used to construct $M_{\{p\}}$. The result trivially follows from Example 1.12 and the fact that
$\operatorname{QR}(\mathbb{F}_{p})=\left\{\,\alpha^{2\ell}\,\,\middle|\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}.$
Note that, unlike in Example 2.5, the ordering of rows indexed using the construction above are not necessarily the ordering as a subsequence of $(1,\,\ldots,\,p-1)$. Proposition 2.8 also implies that other than $\mathbbm{1}$, (p1)/2(p-1)/2 vectors in $\Phi_{p}$ have entries equal to $\zeta$ raised to some permutation of the set of the $\operatorname{QR}$ and (p1)/2(p-1)/2 vectors entries which have $\operatorname{NR}$ as the exponents.
We have one final lemma. As this will also be used to prove results about Paley ETFs $\Phi_{q}$ with $q$ a prime power, we state it in more generality than what is needed for this section.
Lemma 2.9.
Let $q=p^{s}>3$ be a prime power with $q\equiv 3\bmod 4$. Let $X$ be a set of cardinality $q$. Assume that $G$ acts transitively and $2$-homogeneously but not $2$-transitively on $X$. Assume further that
$\tilde{G}=\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}$
is a subgroup of $G$, where
$\operatorname{A\Gamma L}(\mathbb{F}_{q})=\left\{\,f:\mathbb{F}_{q}\rightarrow\mathbb{F}_{q}\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\mathbb{F}_{q}^{\times},\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}.$
Then $G=\tilde{G}$.
Proof.
Theorem 1 of [Kan72] yields that any group which acts $2$-homogeneously but not $2$-transitively on a set of size at least $4$ must be a subgroup of $\operatorname{A\Gamma L}(\mathbb{F}_{q})$. Note that since $q\equiv 3\bmod 4$, $s$ and the cardinality of $\operatorname{QR}(\mathbb{F}_{q})$ must also both be odd, i.e., $\tilde{G}$ has odd order. Also, since $\operatorname{QR}(\mathbb{F}_{q})$ is index $2$ in $\mathbb{F}_{q}^{\times}$, $G$ is a maximal subgroup of $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ of odd order. It follows from the proof of Proposition 3.1 in [Kan69b] that a symmetry group which is $2$-homogeneous but not $2$-transitive must be of odd order. (Alternatively, one may note that $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ acts $2$-transitively on $\mathbb{F}_{q}$ and that $\tilde{G}$ is a maximal proper subgroup.) Hence, $G=\tilde{G}$, as desired. ∎
Note that when $q=p$, a prime, $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ is denoted $\operatorname{AGL}(\mathbb{F}_{p})$ and is parameterized by $\mathbb{F}_{p}$ and $\mathbb{F}_{p}^{\times}$ since there are no nontrivial field automorphisms.
Theorem 2.10.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$. Define $M=M_{\{p\}}$ and $T=T_{(p-1)/2}$. Then
$G:=\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}$
$\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{(p-1)/2}\cong\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,b\in\mathbb{F}_{p},\,\,m\in\operatorname{QR}(\mathbb{F}_{p})\,\right\}$
is the group of vector automorphisms of the Paley ETF $\Phi_{p}$. The vector and line symmetry groups are $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$. The symmetry groups act $2$-homogeneously but not $2$-transitively on the vectors, respectively the lines.
A note of warning for those accustomed to thinking of an affine group as a translation group semidirect product a scaling group, in particular those with a background in wavelet theory: The group of matrices $\{T^{\ell}\}$ which translate the entries of the vectors act like dilations in the correspondence to $\operatorname{AGL}(\mathbb{F}_{p})$.
Proof.
We first show that $\{M^{k}T^{\ell}\}$ is isomorphic to the given subgroup of $\operatorname{AGL}(\mathbb{F}_{p})$ and that $\{M^{k}T^{\ell}\}\leq\operatorname*{Sym}_{v}(\Phi_{p})$. To that end, we note that the order of $M$ is $p$ and the order of $T$ is (p1)/2(p-1)/2. This also means that raising $M$ to elements of $\mathbb{F}_{p}$ is well-defined when interpreted mod $p$. Further, for $\alpha$ the primitive element used to define $M$, $k\in\{0,\,\ldots,\,p-1\}$, and $\ell\in\{0,\,\ldots,\,(p-3)/2\}$, we have by direct computation on the entries of the matrices
$T^{\ell}M^{k}T^{-\ell}=M^{k\alpha^{-2\ell}}.$ (1)
Clearly, $T^{\ell}\mathbbm{1}=\mathbbm{1}$ for any $\ell\in\{0,\,\ldots,\,(p-3)/2\}$. Thus, (1) yields for any frame vector $\varphi_{\alpha^{2k}}$ parameterized by a quadratic residue $\alpha^{2k}$,
$T^{\ell}\varphi_{\alpha^{2k}}=T^{\ell}M^{\alpha^{2k}}\mathbbm{1}=M^{\alpha^{2(k-\ell)}}T^{\ell}\mathbbm{1}=M^{\alpha^{2(k-\ell)}}\mathbbm{1}=\varphi_{\alpha^{2(k-\ell)}}.$
That is, $T^{\ell}$ maps vectors parameterized by quadratic residues to other vectors parameterized by quadratic residues cyclically. A similar calculation shows that left-multiplying by $T$ cyclically permutes the vectors parameterized by the quadratic non-residues. To be clear: $T$ acts on $\mathbb{C}^{(p-1)/2}$ by cyclically permuting the entries of vectors. When applied to $\Phi_{p}$, this cyclic permutation of the rows induces a cyclic permutation of the columns indexed by the quadratic residues and a cyclic permutation of the columns indexed by the quadratic non-residues. So, $\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}$ is a subgroup of the vector automorphisms. The commutation relation in (1) also yields
$\left\{\,M^{k}T^{\ell}\,\,\middle|\,\,k\in\{0,\,\ldots,\,p-1\},\,\,\ell\in\{0,\,\ldots,\,(p-3)/2\}\,\right\}\cong\mathbb{Z}_{p}\rtimes\mathbb{Z}_{(p-1)/2},$
with corresponding group operation
[…]
which may be rewritten as
[…]
which is isomorphic to
$G:=\left\{\,f\in\operatorname{AGL}(\mathbb{F}_{p})\,\,\middle|\,\,f(x)=mx+b,\,\,b\in\mathbb{F}_{p},\,\,m\in\operatorname{QR}(\mathbb{F}_{p})\,\right\}.$
Now we would like to show that $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$ and that the groups act $2$-homogeneously but not $2$-transitively on the vectors/lines.
Note that Proposition 2.8 implies that the vector symmetry group $\operatorname*{Sym}_{v}(\Phi_{p})$ and thus also the line symmetry group $\operatorname*{Sym}_{\ell}(\Phi_{p})$ act transitively since the vectors in $\Phi_{p}$ are the orbit of $\mathbbm{1}$ under $\{M^{k}\}$. We further have that $\{M^{k}T^{\ell}\}$ acts $2$-homogeneously on the vectors as follows. Let $x\neq y$ and $z\neq w$ be in $\mathbb{F}_{p}$. Since $p\equiv 3\bmod 4$, $\operatorname{NR}=-\operatorname{QR}$. In particular either xyx-y or yxy-x is a quadratic residue, and the same is true for $z$ and $w$. Thus, without loss of generality, we pick an ordering so that $x-y,z-w\in\operatorname{QR}$. Then $M^{-y}\varphi_{y}=M^{-w}\varphi_{w}=\mathbbm{1}$, and $M^{-y}\varphi_{x}$ and $M^{-w}\varphi_{z}$ are both columns indexed by quadratic residues. We can then apply a translation $T^{\ell}$ as needed, which fixes $M^{-y}\varphi_{y}=M^{-w}\varphi_{w}=\mathbbm{1}$ and maps $M^{-w}\varphi_{z}$ to $M^{-y}\varphi_{x}$. Thus, $M^{y}T^{-\ell}M^{-w}$ maps $\{\varphi_{z},\varphi_{w}\}$ to $\{\varphi_{x},\varphi_{y}\}$.
We know from [IM24] that $\operatorname*{Sym}_{\ell}(\Phi_{p})$ is not $2$-transitive and thus neither is $\operatorname*{Sym}_{v}(\Phi_{p})$. It thus follows from Lemma 2.9 that $\operatorname*{Sym}_{v}(\Phi_{p})=\operatorname*{Sym}_{\ell}(\Phi_{p})\cong G$. ∎
This theorem (and the more general result Theorem 2.19) should not be incredibly surprising, given that Paley designs are the only $2$-homogeneous but not $2$-transitive Hadamard designs [Kan69b]. Note, however, that the Paley Hadamard designs built from $\mathbb{F}_{11}$ and $\mathbb{F}_{7}$ are $2$-transitive [Tod33, Kan69a], differentiating this result from the similar result for Paley ETFs. As we will see in Theorem 3.5, Paley skew-conference matrices, which are also built using quadratic residues, also yield an infinite class of ETFs which are $3$-homogeneous but not $3$-transitive.
2.3 Alternate Proof of Theorem 2.10
In this section, we will provide an alternate proof characterizing the symmetry groups of $\Phi_{p}$ which filters through the inner products, respectively, triple products (i.e., uses Theorem 1.14). This has the benefit that some of the techniques might be useful in the analysis of symmetry groups of other ETFs but the drawback that the explicit unitary matrices which are vector, respectively, line automorphisms are not constructed. We conclude with a short example of using the approach on another class of ETFs (Gabor-Steiner). In order to use Theorem 1.14, we need to characterize the inner products and triple products of the Paley ETF vectors.
Lemma 2.11.
Let $p>3$ be a prime with $p\equiv 3\bmod 4$ and consider the Paley ETF $\Phi_{p}$. Further let $\zeta$ be the primitive $p$th root of unity and $\alpha^{2}$ the generator of $\operatorname{QR}(\mathbb{F}_{p})$ used to produce $\Phi_{p}$. Finally, set
$a=\sum_{\ell=0}^{(p-3)/2}\zeta^{\alpha^{2\ell}}.$
Then
$\operatorname*{TP}(j,k,\ell)=\left\{\begin{array}[]{lr}a^{3};&j-k,k-\ell,\ell-j\in\operatorname{QR}\\[2.0pt] \left|a\right|^{2}a;&\textrm{ two of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \left|a\right|^{2}\overline{a};&\textrm{ one of $j-k,k-\ell,\ell-j\in\operatorname{QR}$}\\[2.0pt] \overline{a}^{3};&\ j-k,k-\ell,\ell-j\in\operatorname{NR}\end{array}\right$.
Proof.
First, we note that
$\left\langle\varphi_{j},\varphi_{k}\right\rangle=\sum_{\ell=0}^{(p-3)/2}\zeta^{j\alpha^{2\ell}}\zeta^{-k\alpha^{2\ell}}=\sum_{\ell=0}^{(p-3)/2}\zeta^{(j-k)\alpha^{2\ell}}=\left\{\begin{array}[]{lr}(p-1)/2&j=k\\ a&j-k\in\operatorname{QR}\\ \overline{a}&j-k\in\operatorname{NR}\end{array}\right$.
since $(j-k)\alpha^{2\ell}\in\operatorname{QR}$ if and only if $j-k\in\operatorname{QR}$ and also $\operatorname{NR}=-\operatorname{QR}$. So, for distinct $j,k,\ell$, the desired result immediately follows. ∎
These triple products are distinct.
Lemma 2.12.
Let $a$ be as in Lemma 2.11. Then $a^{3}$, $\left|a\right|^{2}a$, $\left|a\right|^{2}\overline{a}$, and $\overline{a}^{3}$ are pairwise distinct.
Proof.
At least one pair of the numbers in the above list is equal if and only if $a\in\{\overline{a},-\overline{a},\overline{a}\eta,\overline{a}\eta^{2}\}$ for $\eta$ a primitive third root of unity. Let $\omega$ be any sixth root of unity. Note that $a$ and $\overline{a}$ are sums of (p1)/2(p-1)/2 distinct primitive $p$th roots of unity (indexed by the quadratic residues and quadratic non-residues, respectively). Thus, $a-\overline{a}\omega$ is a rational linear combination of p1p-1 distinct $6p$th roots of unity, where the smallest order of the ratio of any two of the roots of unity is $p$ (since $6$ and $p$ are relatively prime). Theorem 2.6 in [Len79] (see also [CJ76]) yields that a rational linear combination of roots of unity, where the smallest order of the ratio of any two of the roots of unity is a prime $p$, can never vanish if there are fewer than $p$ terms. Thus, $a-\overline{a}\omega$ cannot be 0, yielding that $a^{3}$, $\left|a\right|^{2}a$, $\left|a\right|^{2}\overline{a}$ are four distinct numbers. ∎
We are now ready to give an alternate proof of Theorem 2.10.
Alternate Proof of Theorem 2.10.
It follows from Lemma 2.11 that the value of the inner product of two distinct vectors in $\Phi_{p}$ depends solely on whether the difference of the indices is a quadratic residue. In particular, if we cyclically shift all of the indices by $\ell\in\mathbb{F}_{p}$, none of the inner products or triple products change (as $(j-\ell)-(k-\ell)=j$-$\ell$). This means that $\operatorname*{Sym}_{v}(\Phi_{p})$ acts transitively. Also, note that multiplying the indices by a quadratic residue does not change the inner products or triple products since $\operatorname{QR}$ is a subgroup of $\mathbb{F}_{p}^{\times}$ (i.e., $\alpha^{2\ell}j-\alpha^{2\ell}k=\alpha^{2\ell}(j-k)\in\operatorname{QR}\Leftrightarrow j-k\in\operatorname{QR}$).
Let $x\neq y$ and $z\neq w$ be in $\mathbb{F}_{p}$. We may choose without loss of generality an ordering such that $x-y,z-w\in\operatorname{QR}$. Then there is a quadratic residue $\alpha^{2\ell}$ that scales xyx-y to zwz-w. So, we apply the following sequence of permutations:
$(x,y)\,\rightarrow\,(x-y,y-y)=(x-y,0)\,\rightarrow\,(\alpha^{2\ell}(x-y),\alpha^{2\ell}(0))=(z-w,0)\,\rightarrow\,((z-w)+w,0+w)=(z,w).$
Hence, $\operatorname*{Sym}_{v}(\Phi_{p})$ acts $2$-homogeneously. We know from [IM24] that $\operatorname*{Sym}_{\ell}(\Phi_{p})$ is not $2$-transitive and thus neither is $\operatorname*{Sym}_{v}(\Phi_{p})$. The desired result now follows from Lemma 2.9. ∎
As an another example of this technique, we consider the Gabor-Steiner ETFs [BK19]. (See also related constructions in [BE10a, BE10b, IJM20, FJM+19].) Namely, each finite abelian group $G$ of odd order $m$ generates a so-called Gabor-Steiner ETF $\Phi_{G}$ of $m^{2}$ vectors in $\mathbb{C}^{m(m-1)/2}$. The vectors of $\Phi_{G}$ are the orbit of a projective unitary representation of $G\times G$ and are thus parameterized by $G\times G$. For $G=(\mathbb{F}_{p},+)$ with $p$ an odd prime and $\zeta$ a primitive $p$th root of unity, Lemma 5.2 of [BK19] yields that the triple products of the vectors are
$\operatorname*{TP}((k,\kappa),(\tilde{k},\tilde{\kappa}),(\widehat{k},\widehat{\kappa}))=\zeta^{[(k\widehat{\kappa}-\widehat{k}\kappa)+(\tilde{k}\kappa-k\tilde{\kappa})+(\widehat{k}\tilde{\kappa}-\tilde{k}\widehat{\kappa})](p+1)/2}.$
One can see from inspection (i.e., just writing out the coordinates and simplifying) that for $\operatorname{Sp}(2s,p)\subset\mathbb{F}_{p}^{2s\times 2s}$ the group of $2s\times 2s$ symplectic matrices,
$\operatorname{ASp}(2,p)=\left\{\,f:\mathbb{F}_{p}^{2}\rightarrow\mathbb{F}_{p}^{2}\,\,\middle|\,\,f(x)=Mx+b,\,\,M\in\operatorname{Sp}(2,p),\,\,b\in\mathbb{F}_{p}^{2}\,\right\}\leq\operatorname*{Sym}_{\ell}(\Phi_{(\mathbb{F}_{p},+)})$
[…]
2.4 Paley ETFs over Fields of Prime Power Order
We seek to prove a generalization of Theorem 2.10 which holds for any Paley ETF, not just those built over a field of prime order. In the proof of Theorem 2.10, we were easily able to relate the multiplicative structure of the quadratic residues, which was used to index the rows, with the exponents in the entries of the vectors. When dealing with Paley ETFs over $\mathbb{F}_{p^{s}}$ for $s>1$, we will need to connect the multiplicative structure of $\operatorname{QR}$ in $\mathbb{F}_{p^{s}}^{\times}$ with the additive structure of the vector space $\mathbb{F}_{p}^{s}$ in a way that works regardless of the minimal polynomial used to represent $\mathbb{F}_{p^{s}}$ as a Galois extension of $\mathbb{F}_{p}$.
We need to define unitary operators supplementing those in Definition 2.7, which we will prove in Theorem 2.19 form the line automorphism group. In particular, we need to define an operator that corresponds to the Galois group action in $\operatorname{A\Gamma L}(\mathbb{F}_{q})$ (cf. Lemma 2.9).
Definition 2.13.
Let $q=p^{s}>3$ be a prime power with $q\equiv 3\bmod 4$. Let $\Phi_{q}$ be the Paley ETF with rows indexed by $\left\{\,\alpha^{2k}\,\,\middle|\,\,k\in\{0,\ldots,(q-3)/2\}\,\right\}$ for $\alpha$ a primitive element. The order-$q$ Paley modulation group consists of $M_{\{q\}}^{\{b\}}\in\operatorname{U}((q-1)/2)$ the diagonal matrix with diagonal elements the vector in $\Phi_{q}$ indexed by $b$. The Galois permutation $\Pi_{q}\in\operatorname{U}((q-1)/2)$ maps coordinate $\ell$ to coordinate $p\ell\bmod(q-1)/2$.
As before, when the parameters are clear from context, we will omit the subscripts. Note that
$\left\{\,M_{\{q\}}^{\{b\}}\,\,\middle|\,\,b\in\mathbb{F}_{q}\,\right\}$
forms a group isomorphic to $(\mathbb{F}_{q},+)$. This follows from the character table structure of the Fourier matrices (Example 1.12). Galois theory also yields that $\Pi_{q}$ is well-defined and order $s$. When $q=p$ is a prime, $S$ is the $1\times 1$ identity over $\mathbb{F}_{p}$ and $\Pi$ the $(p-1)/2\times(p-1)/2$ identity over $\mathbb{C}$.
Example 2.14.
We define the operators of note for $\Phi_{27}$. $T_{13}$ is the $13\times 13$ cyclic translation matrix that implements (via left multiplication) the row permutation $(0\,1\,2\,3\,\cdots\,12)$. $\Pi_{27}$ is the $13\times 13$ permutation matrix that implements (via left multiplication) the row permutation (0)(1 3 9)(2 6 5)(4 12 10)(7 8 11)(0)(1\,3\,9)(2\,6\,5)(4\,12\,10)(7\,8\,11). Note that 0 is fixed because it corresponds to $1=\alpha^{0}$, i.e., the lone quadratic residue in the base field $\mathbb{F}_{3}$. Finally, for $\zeta$ the primitive third root of unity used to define $\Phi_{27}$, the modulation group is
$\{I\}\cup\left\{\,\operatorname*{diag}\left(\left(\zeta^{\left\langle\alpha^{2k},\alpha^{\ell}\right\rangle_{3}}\right)_{k\in\{0,\ldots,12\}}\right)\,\,\middle|\,\,\ell\in\{0,\ldots,25\}\,\right\}.$
We would like to show that $T_{(q-1)/2}$ and $\Pi_{q}$ are indeed vector automorphisms of $\Phi_{q}$. One might hope that once we have fixed a primitive element $\alpha$ with corresponding companion matrix $C$, partitioning the vectors of $\Phi_{q}$ into sets parameterized by 0, $\operatorname{QR}$ (i.e., $C^{\ell}e_{1}$ for $\ell$ even), and $\operatorname{NR}$ (i.e., $C^{\ell}e_{1}$ for $\ell$ odd) would allow us to do analysis as in the proof of Theorem 2.10. However, the actual partitioning will rely on the image of these sets under a special invertible map. Proposition 2.16 and Lemma 2.17 concern the existence of an invertible map with the desired properties. First, we note that in the fields pertinent for Paley ETFs, there will always be a normal basis consisting of quadratic residues.
Lemma 2.15.
Let $q\equiv 3\bmod 4$ where $q=p^{s}$ for $p$ a prime. Then, there exists a normal basis of $\mathbb{F}_{q}$ which is the orbit of an $x\in\operatorname{QR}$, i.e., $\left\{\,x^{p^{i}}\,\,\middle|\,\,i\in\{0,1,\ldots,s-1\}\,\right\}\subset\operatorname{QR}$.
Proof.
By the normal basis theorem (see, e.g., [Art59]), $\mathbb{F}_{q}$ must have at least one normal basis, which clearly cannot be the orbit of 0. Consider a normal basis $\left\{\,x^{p^{i}}\,\,\middle|\,\,i\in\{0,1,\ldots,s-1\}\,\right\}$ which is the orbit of a quadratic non-residue, $x$. Since $q\equiv 3\bmod 4$, x-x is a quadratic residue. As $\operatorname{QR}$ forms a subgroup of $\mathbb{F}_{q}^{\times}$, the orbit of x-x under the Galois group must only consist of quadratic residues. Also, since for each $i\in\{0,1,\ldots,s-1\}$, $(-x)^{p^{i}}=(-1)^{p^{i}}x^{p^{i}}=-x^{p^{i}}$, the orbit of x-x under the Galois group must be linearly independent and also form a basis for $\mathbb{F}_{q}$. ∎
Proposition 2.16.
Let $M\in\mathbb{F}^{s\times s}$ be similar to a companion matrix of a separable irreducible polynomial $\pi\in\mathbb{F}[X]$. Further, let $(f_{k})_{k=0}^{s-2}$ be s1s-1 linear functionals on $\mathbb{F}^{s\times s}$. Then there exists an invertible symmetric matrix $S\in\mathbb{F}^{s\times s}$ such that the following hold
  1. 1.
    $SM=M^{\top}S$ and
  2. 2.
    $f_{k}(S)=0$ for all $k\in\{0,\ldots,s-2\}$.
Proof.
Since $M$ is similar to a companion matrix, Theorems 1 & 2 of [TZ59] yield that there is a symmetric solution to (1.) and that all solutions to (1.) are symmetric. The proof of these theorems in [TZ59] shows that the space of solutions to (1.) is
$\left\{\,S_{0}\left(a_{0}I_{s}+a_{1}M+\ldots+a_{s-1}M^{s-1}\right)\,\,\middle|\,\,a_{0},a_{1},\ldots,a_{s-1}\in\mathbb{F}\,\right\}$
for $S_{0}$ an invertible matrix which is also a solution to (1.). First note that this structure means that the solutions to (1.) form an $s$-dimensional vector space over $\mathbb{F}$. Combining this with the fact that (2.) is a system of s1s-1 homogeneous linear equations yields that the solution set to the system of (1.) and (2.) must have dimension at least s(s1)=1s-(s-1)=1. We now need to show that there is an invertible element of the solution set. We will use the additional restrictions on the minimal polynomial of $M$ to show that every non-zero solution to (1.) is invertible. This is equivalent to showing that, other than the zero matrix, the vector space of degree at-most s1s-1 polynomials in $\mathbb{F}[X]$ evaluated at $M$ consists of only invertible matrices. Consider a field extension $\mathbb{K}$ of $\mathbb{F}$ that contains the roots of $\pi$. Since $\pi$ is separable over $\mathbb{K}$, $M$ is similar to a diagonal Jordan block. Further, since $\pi$ is irreducible, no non-zero polynomial of degree <s<s with coefficients in $\mathbb{F}$ evaluated on the roots of $\pi$ are zero; thus, no such polynomial evaluated on the diagonalized $M$ will yield any zeros on the diagonal. Hence, the evaluation is invertible over $\mathbb{K}$ and thus over $\mathbb{F}$, as desired. ∎
The above proposition helps to characterize the exponents in the entries of the Paley ETFs in a way that will allow us to analyze the actions of $T_{(q-1)/2}$ and $\Pi_{q}$.
Lemma 2.17.
Let $q=p^{s}\equiv 3\bmod 4$ be a prime power. Let $\pi\in\mathbb{F}_{p}[X]$ be the polynomial used to generate $\mathbb{F}_{q}$, i.e., $\mathbb{F}_{q}\cong\mathbb{F}_{p}[X]/(\pi(X))$. If $C\in\mathbb{F}_{p}^{s\times s}$ is the companion matrix of $\pi$ and $e_{1}\in\mathbb{F}_{p}^{s}$ is the first standard basis vector, there exists an invertible symmetric $S\in\mathbb{F}_{p}^{s\times s}$ such that
  1. 1.
    $\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2(k+n)}e_{1},SC^{2(\ell-n)}e_{1}\right\rangle_{p}$ for all $k,\ell,n\in\{0,\ldots,(q-3)/2\}$ and
  2. 2.
    $\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2pk}e_{1},SC^{2p\ell}e_{1}\right\rangle_{p}$ for all $k,\ell\in\{0,\ldots,(q-3)/2\}$.
Proof.
Since $q\equiv 3\bmod 4$, Lemma 2.15 guarantees the existence of a basis consisting of the Galois orbit of a quadratic residue. Let $x=C^{2b}e_{1}$ denote one such quadratic residue which generates a normal basis. Further, since $C$ is a companion matrix for a separable irreducible polynomial, Proposition 2.16 shows the existence of an invertible symmetric $S\in\mathbb{F}^{s\times s}$ which solves the system
  1. 1’.
    $SC=C^{\top}S$ and
  2. 2’.
    $e_{1}^{\top}S\left(C^{2bp^{k}}-C^{2bp^{k+1}}\right)e_{1}=0$ for all $k\in\{0,\ldots,s-2\}$.
We seek to show that this $S$ also solves (1.) and (2.). First, (1’.) implies that for all $k,\ell,n\in\{0,\ldots,(q-3)/2\}$
$\left(C^{2k}\right)^{\top}SC^{2\ell}$ $=\left(C^{2(k+n)}\right)^{\top}\left(C^{-2n}\right)^{\top}SC^{2\ell}$
$=\left(C^{2(k+n)}\right)^{\top}S\left(C^{-2n}\right)C^{2\ell}$
$=\left(C^{2(k+n)}\right)^{\top}SC^{2(\ell-n)}$
$\hskip-55.0pt\Rightarrow e_{1}^{\top}\left(C^{2k}\right)^{\top}SC^{2\ell}e_{1}=e_{1}^{\top}\left(C^{2(k+n)}\right)^{\top}SC^{2(\ell-n)}e_{1}$
$\hskip-55.0pt\Rightarrow\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2(k+n)}e_{1},SC^{2(\ell-n)}e_{1}\right\rangle_{p};$
that is, $S$ solves (1.). Now we use (1.’) and (2.’) to show that (2.) holds. First note that (2.’) yields
$e_{1}^{\top}SC^{2b}e_{1}=e_{1}^{\top}SC^{2bp}e_{1}=\ldots=e_{1}^{\top}SC^{2bp^{s-1}}e_{1}.$
Let $k,\ell\in\{0,\ldots,(q-3)/2\}$ and set $y=C^{2(k+\ell)}e_{1}$. Since $\left\{\,x^{p^{n}}=C^{2bp^{n}}e_{1}\,\,\middle|\,\,n\in\{0,\ldots,s-1\}\,\right\}$ is a basis for $\mathbb{F}_{q}$ as a vector space over $\mathbb{F}_{p}$, there exists $c_{n}\in\mathbb{F}_{p}$ such that $y=\sum_{n=0}^{s-1}c_{n}x^{p^{n}}$. As the Frobenius automorphism fixes $\mathbb{F}_{p}$, we have $y^{p}=\sum_{n=0}^{s-1}c_{n}x^{p^{n+1}}$. Rewriting this using vector space notation, we have
$C^{2(k+\ell)}e_{1}=\sum_{n=0}^{s-1}c_{n}C^{2bp^{n}}e_{1}\quad\textrm{and}\quad C^{2p(k+\ell)}e_{1}=\sum_{n=0}^{s-1}c_{n}C^{2bp^{n+1}}e_{1}.$
Now, using (1.’) and (2.’), we have that
$\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}$ $=e_{1}^{\top}\left(C^{2k}\right)^{\top}SC^{2\ell}e_{1}$
$=e_{1}^{\top}SC^{2(k+\ell)}e_{1}$
$=e_{1}^{\top}S\sum_{n=0}^{s-1}c_{n}C^{2bp^{n}}e_{1}$
$=\sum_{n=0}^{s-1}c_{n}e_{1}^{\top}SC^{2bp^{n}}e_{1}$
$=\sum_{n=0}^{s-1}c_{n}e_{1}^{\top}SC^{2bp^{n+1}}e_{1}$
$=e_{1}^{\top}SC^{2p(k+\ell)}e_{1}$
$=\left\langle C^{2pk}e_{1},SC^{2p\ell}e_{1}\right\rangle_{p},$
as desired. ∎
Example 2.18.
Continuing Examples 2.22.3, and 2.6, let
$S=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&1\end{pmatrix}$.
Note, e.g., that
$C^{\top}S=SC=\begin{pmatrix}0&1&0\\ 1&0&1\\ 0&1&2\end{pmatrix}\quad\textrm{and}\quad\left\langle C^{2}e_{1},SC^{4}e_{1}\right\rangle_{3}=\left\langle C^{6}e_{1},SC^{12}e_{1}\right\rangle_{3}=1.$
This $S$ is indeed a solution to the system in Lemma 2.17.
We are now ready to present the main theorem of this section.
Theorem 2.19.
Let $q=p^{s}>3$ be a prime power with $q\equiv 3\bmod 4$. Define $M^{\{b\}}=M^{\{b\}}_{\{q\}}$, $T=T_{(q-1)/2}$, and $\Pi=\Pi_{q}$. Then
$G:=\left\{\,M^{\{b\}}T^{\ell}\Pi^{k}\,\,\middle|\,\,b\in\mathbb{F}_{q},\,\,\ell\in\{0,\,\ldots,\,(q-3)/2\},\,\,k\in\{0,\,\ldots,\,s-1\}\,\right\}$
$\cong\left(\left(\mathbb{Z}_{p}\right)^{s}\rtimes\mathbb{Z}_{(q-1)/2}\right)\rtimes\mathbb{Z}_{s}\cong\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}.$
is the group of vector automorphisms of the Paley ETF $\Phi_{q}$. The vector and line symmetry groups are $\operatorname*{Sym}_{v}(\Phi_{q})=\operatorname*{Sym}_{\ell}(\Phi_{q})\cong G$. The symmetry groups act $2$-homogeneously but not $2$-transitively on the vectors and the lines.
Proof.
We first claim that the modulation group, the cyclic translation, and the Galois permutation are all vector automorphisms of $\Phi_{q}$ and that $\{M^{\{b\}}T^{\ell}\Pi^{k}\}$ is isomorphic to the desired subgroup of $\operatorname{A\Gamma L}(\mathbb{F}_{q})$. As in Proposition 2.8, the vectors of $\Phi_{q}$ are trivially the orbit of $\varphi_{0}=\mathbbm{1}$ under the modulation group. Note that this also implies that $\operatorname*{Sym}_{v}(\Phi_{q})\leq\operatorname*{Sym}_{\ell}(\Phi_{q})$ act transitively on the vectors and lines, respectively.
Let $S$ be a matrix defined as in Lemma 2.17 and $e_{1}\in\mathbb{F}_{p}^{s}$ be the first standard basis vector. Since $S$ is linear and invertible, we may partition the indices of the vectors of $\Phi_{q}$ as 0, the image of the quadratic residues under $S$, i.e.,
$S(\operatorname{QR})=\left\{\,SC^{2\ell}e_{1}\,\,\middle|\,\,\ell\in\{0,\ldots,(q-3)/2\}\,\right\},$
and the image of the quadratic non-residues under $S$, i.e.,
$S(\operatorname{NR})=S(-\operatorname{QR})=-S(\operatorname{QR}).$
We trivially have that $T\mathbbm{1}=\mathbbm{1}$. We claim that left-multiplying by $T$ has the effect of cyclically permuting the frame vectors indexed by $S(\operatorname{QR})$ and also cyclically permuting the frame vectors indexed by $S(\operatorname{NR})$. Let $\zeta$ be the primitive $p$th root of unity used to generate $\Phi_{q}$. Note that the $k$th entry of the vector indexed by $SC^{2\ell}e_{1}$ is $\zeta^{\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}}$. It follows from Lemma 2.17 that $\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2(k+1)}e_{1},SC^{2(\ell-1)}e_{1}\right\rangle_{p}$ for all $k,\ell\in\{0,\ldots,(q-3)/2\}$. Thus,
$T\varphi_{SC^{2\ell}e_{1}}=\varphi_{SC^{2(\ell-1)}e_{1}}\quad\textrm{for all}\quad\ell\in\{0,\ldots,(q-3)/2\};$
i.e., $T$ cyclically permutes the vectors indexed by $S(\operatorname{QR})$. Since $S$ is linear and $\operatorname{NR}=-\operatorname{QR}$, we also get that $\left\langle C^{2k}e_{1},-SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2(k+1)}e_{1},-SC^{2(\ell-1)}e_{1}\right\rangle_{p}$ for all $k,\ell\in\{0,\ldots,(q-3)/2\}$, i.e., $T$ cyclically permutes the vectors indexed by $S(\operatorname{NR})$. Since each $M^{\{b\}}$ is the diagonal matrix with diagonal entries $\varphi_{b}$, this means that $T^{\ell}M^{\{b\}}T^{-\ell}=M^{\{SC^{-2\ell}S^{-1}b\}}$.
Similarly, note that $\Pi\mathbbm{1}=\mathbbm{1}$. We will show that left multiplying by $\Pi$ has the effect of moving a vector indexed by $Sx$ for some $x\in\mathbb{F}_{q}$ to $Sx^{p}$. Since the Frobenius automorphism maps quadratic residues to quadratic residues and quadratic non-residues to quadratic non-residues, we will again analyze $S(\operatorname{QR})$ first. It follows from Lemma 2.17 that $\left\langle C^{2k}e_{1},SC^{2\ell}e_{1}\right\rangle_{p}=\left\langle C^{2pk}e_{1},SC^{2p\ell}e_{1}\right\rangle_{p}$ for all $k,\ell\in\{0,\ldots,(q-3)/2\}$. Thus,
$\Pi\varphi_{SC^{2\ell}e_{1}}=\varphi_{SC^{2p\ell}e_{1}}\quad\textrm{for all}\quad\ell\in\{0,\ldots,(q-3)/2\};$
i.e., $\Pi$ maps $Sx$ to $Sx^{p}$ for any $x\in\operatorname{QR}$. Similarly, $\Pi$ maps $Sx$ to $Sx^{p}$ for any $x\in\operatorname{NR}$. Using these calculations, we may compute the remaining commutator relations: $\Pi^{k}M^{\{b\}}\Pi^{-k}=M^{\{S(S^{-1}b)^{p^{k}}\}}$ and $\Pi^{k}T^{\ell}\Pi^{-k}=T^{p^{k}\ell}$.
We would like to show that
$\left\{\,M^{\{b\}}T^{\ell}\Pi^{k}\,\,\middle|\,\,b\in\mathbb{F}_{q},\,\,\ell\in\{0,\,\ldots,\,(q-3)/2\},\,\,k\in\{0,\,\ldots,\,s-1\}\,\right\}$
has the desired structure. Tracking the permutation induced on the preimage of the vector indices under $S$ and using the commutation relations above, we see that the product of two operators of the form $M^{\{b\}}T^{\ell}\Pi^{k}$ yields the product
[…]
which is isomorphic to
[…]
which in turn is isomorphic to
$\left(\left(\mathbb{Z}_{p}\right)^{s}\rtimes\mathbb{Z}_{(q-1)/2}\right)\rtimes\mathbb{Z}_{s}\cong\left\{\,f\in\operatorname{A\Gamma L}(\mathbb{F}_{q})\,\,\middle|\,\,f(x)=mx^{\sigma}+b,\,\,b\in\mathbb{F}_{q},\,\,m\in\operatorname{QR}(\mathbb{F}_{q}),\,\,\sigma\in\operatorname{Aut}(\mathbb{F}_{q})\,\right\}.$
Thus, $G\leq\operatorname*{Sym}_{v}(\Phi_{q})\leq\operatorname*{Sym}_{\ell}(\Phi_{q})$.
The remainder of the proof is almost exactly like the end of the proof of Theorem 2.10. Namely, consider $x\neq y$ and $z\neq w$ in $\mathbb{F}_{q}$. Then either $S^{-1}(x-y)$ or $S^{-1}(y-x)$ is a quadratic residue, similarly for $(z,w)$. Thus, without loss of generality, we pick an ordering so that $S^{-1}(x-y),S^{-1}(z-w)\in\operatorname{QR}$. Then $M^{\{-y\}}\varphi_{y}=M^{\{-w\}}\varphi_{w}=\mathbbm{1}$, and $M^{\{-y\}}\varphi_{x}$ and $M^{\{-w\}}\varphi_{z}$ are both columns indexed by $S(\operatorname{QR})$. We can then apply a translation $T^{\ell}$ as needed, which fixes $\mathbbm{1}$ and maps $M^{\{-w\}}\varphi_{z}$ to $M^{\{-y\}}\varphi_{x}$. Thus, $M^{\{y\}}T^{-\ell}M^{-\{w\}}$ maps $\{\varphi_{z},\varphi_{w}\}$ to $\{\varphi_{x},\varphi_{y}\}$. Hence, $\operatorname*{Sym}_{v}(\Phi_{q})\leq\operatorname*{Sym}_{\ell}(\Phi_{q})$ act $2$-homogeneously on the vectors and lines, respectively.
We know from [IM24] that $\operatorname*{Sym}_{\ell}(\Phi_{q})$ is not $2$-transitive and thus neither is $\operatorname*{Sym}_{v}(\Phi_{q})$. It thus follows from Lemma 2.9 that $\operatorname*{Sym}_{v}(\Phi_{q})=\operatorname*{Sym}_{\ell}(\Phi_{q})\cong G$. ∎
Corollary 2.20.
Let $q>3$ be a prime power with $q\equiv 3\bmod 4$. The bender of the Paley ETF $\Phi_{q}$ is a BIBD.
Proof.
This follows immediately from Theorem 2.19 and Proposition 1.23
Due to a result from Chebotarëv [SL96], the Paley ETFs $\Phi_{p}$ with $p$ prime are full spark, i.e., have spark (p1)/2+1=(p+1)/2(p-1)/2+1=(p+1)/2. Their bender is just trivially the set of all subsets of size $(p+1)/2$, which is a $t$-design for all $1\leq t\leq(p+1)/2$. For a discussion of $p^{s}$ with $s>1$, see Question 4.1.
3 $k$-Homogeneous ETFs
In this section, we characterize $k$-homogeneous ETFs. Throughout this section, we assume $d>1$ as otherwise $d>1$ would need to be added to the statement of almost every result. In particular, this means that the ETFs cannot trivially contain repeated elements.
The following result is generally known, but we include a proof here for completeness.
Lemma 3.1.
The line symmetry group of orthonormal bases and simplex ETFs with $n$ vectors is $S_{n}$. If an ETF is a regular simplex ETF or an orthonormal basis, then the vector symmetry group is also isomorphic to $S_{n}$.
Proof.
Given any two ordered orthonormal bases $U_{1}$ and $U_{2}$ of $\mathbb{C}^{n}$, $U_{2}U_{1}^{-1}$ is a unitary matrix that maps the basis $U_{1}$ to the basis $U_{2}$. Thus, the automorphism group of the vectors and in turn the lines is $S_{n}$. A $(d,n)$-ETF has $n=d+1$ vectors if and only if all of the triple products are negative [FJKM18]. Thus, by Theorem 1.14, any permutation of the lines is in the symmetry group. Finally, if the ETF is a regular simplex ETF, then the inner products are all negative and thus any permutation of the vectors is possible (see, e.g., [CW16]). ∎
We also include the following converse.
Lemma 3.2.
The only ETFs with $3$-transitive line symmetry groups are orthonormal bases and simplices. The only ETFs with $2$-transitive vector symmetry groups are orthonormal bases and regular simplex ETFs.
Proof.
The first result is Theorem II.6 in [Kin19]. The second comes from the fact that if the vector symmetry group is $2$-transitive, then any inner product of distinct vectors is mapped to any other inner product. ∎
Note that this gives an alternate proof that the vector symmetry group of the prime order Paley ETFs $\Phi_{p}$ is not $2$-transitive. We note that the following immediately follows from the definition of homogeneity, but we record it here as we will use it frequently.
Lemma 3.3.
If a group acts $k$-homogeneously on a set of size $n$, then it also acts (nk)(n-k)-homogeneously.
We now have a final lemma before stating and proving the main result of this section.
Lemma 3.4.
If $\Phi$ is an ETF with $n>3$ vectors which has $3$-homogeneous line symmetry group, then up to switching equivalence, the inner products are all 0 ($n=d$), negative ($n=d+1$), or purely imaginary ($n=2d$).
Proof.
Let $\Phi=(\varphi_{j})_{j}$ be arbitrary and let $j,k,\ell$ be indices. Note that
$\overline{\operatorname*{TP}(j,k,\ell)}=\overline{\left\langle\varphi_{j},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{j}\right\rangle}$ $=\left\langle\varphi_{k},\varphi_{j}\right\rangle\left\langle\varphi_{\ell},\varphi_{k}\right\rangle\left\langle\varphi_{j},\varphi_{\ell}\right\rangle=\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\left\langle\varphi_{\ell},\varphi_{k}\right\rangle\left\langle\varphi_{k},\varphi_{j}\right\rangle=\operatorname*{TP}(j,\ell,k).$
Also note that triple products are invariant to cyclic permutation of the indices; i.e.,
$\operatorname*{TP}(j,k,\ell)=\operatorname*{TP}(\ell,j,k)=\operatorname*{TP}(k,\ell,j).$
Thus, for any fixed set of indices $\{j,k,\ell\}$, there are only two possible triple products regardless of order, $\operatorname*{TP}(j,k,\ell)$ and $\overline{\operatorname*{TP}(j,k,\ell)}$. Now assume that $\Phi$ is an ETF with $n>3$ vectors which has a $3$-homogeneous line symmetry group. Since $\Phi$ is an ETF, every triple product (of distinct indices) has modulus $\left|a\right|^{3}$ for some $a\in\mathbb{C}$. And further, since $\Phi$ is $3$-homogeneous and every set of three lines is mapped to every other set of three lines, $\operatorname*{TP}(j,k,\ell)\in\{\operatorname*{TP}(\tilde{j},\tilde{k},\tilde{\ell}),\,\overline{\operatorname*{TP}(\tilde{j},\tilde{k},\tilde{\ell})}\}$ for any distinct triples of distinct indices. That is, every triple product must be of the form $a^{3}$ or $\overline{a^{3}}$. Assume without loss of generality that the norms of the vectors of the ETF are all $1$. An ETF with unit norm vectors such that the spectral norm of
$\begin{pmatrix}1&\left\langle\varphi_{j},\varphi_{k}\right\rangle&\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{k},\varphi_{j}\right\rangle&1&\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{\ell},\varphi_{j}\right\rangle&\left\langle\varphi_{\ell},\varphi_{k}\right\rangle&1\end{pmatrix}$
is constant for all $j\neq k\neq\ell\neq j$ is called $3$-uniform [BP05] (or $3_{c}$-uniform [HS12]). It was shown in [HS12] that the only $3_{c}$-uniform ETFs are, up to switching equivalence, orthonormal bases, simplices, or $(d,2d)$-ETFs with purely imaginary inner products. Thus, we would like to show that $3$-homogeneous ETFs are $3_{c}$-uniform. To that end, we compute the characteristic polynomial of the Gram submatrix above:
$\det\begin{pmatrix}1-\lambda&\left\langle\varphi_{j},\varphi_{k}\right\rangle&\left\langle\varphi_{j},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{k},\varphi_{j}\right\rangle&1-\lambda&\left\langle\varphi_{k},\varphi_{\ell}\right\rangle\\ \left\langle\varphi_{\ell},\varphi_{j}\right\rangle&\left\langle\varphi_{\ell},\varphi_{k}\right\rangle&1-\lambda\end{pmatrix}$
$=(1-\lambda)^{3}-(1-\lambda)\left(\left|\langle\varphi_{j},\varphi_{k}\rangle\right|^{2}+\left|\langle\varphi_{j},\varphi_{\ell}\rangle\right|^{2}+\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|^{2}\right)+\operatorname*{TP}(j,k,\ell)+\overline{\operatorname*{TP}(j,k,\ell)}$
$=(1-\lambda)^{3}-3(1-\lambda)\left|a\right|^{2}+2\operatorname{Re}a^{3},$
which is constant for any distinct $j,k,\ell$. Thus, the spectral norms of any of the $3\times 3$ Gram submatrices are equal to each other as their spectra are all equal. ∎
The concept of $k$-uniformity was introduced to study robustness of frames to $k$ erasures [HP04, BP05]. (Optimal total volume [CC23] of Parseval frames is also a related concept.) Equal-norm tight frames are optimal under certain models to noise and single erasures [GKK01] and are in turn $1$-uniform. ETFs are optimal to two erasures and are $2$-uniform [BP05, SH03, HP04]. ETFs with all zero, all negative, or all purely imaginary triple products are optimal to $3$ erasures and $3$-uniform [BP05, HS12]. Symmetry and $k$-uniformity are related, with $k$-fold symmetry almost implying $k$-uniformity, as outlined below. It is known that a Fourier frame, i.e., a frame formed as an all-column submatrix of a Fourier transform matrix is always an equal-norm tight frame with transitive vector symmetry group containing the group of modulations (similar to Proposition 2.8). More generally, the orbit of a non-zero vector under a finite irreducible unitary representation is always an equal norm tight frame [VW05] with transitive vector symmetry group. There are other results concerning when orbits of a non-zero vector under a projective unitary representation yields an equal-norm tight frame with necessarily transitive line symmetry group (see, e.g., [LPW05, ACGO25]). It is, however, possible for a spanning set to have a transitive vector or line symmetry group and not be a tight frame. For example, if one multiplies the first entry of every vector in a Paley ETF $\Phi_{p}$ by a non-unimodular scalar, the resulting set of vectors form a spanning set with a transitive vector/line symmetry group that contains cyclic permutations induced from the Paley modulation, but the new vectors would no longer be a tight frame / $1$-uniform frame. Any spanning set which has $2$-transitive line symmetry group is an ETF [IM24], and it immediately follows from the definition of $2$-homogeneity that any equal-norm tight frame which has $2$-homogeneous line symmetry group is an ETF, where ETFs are precisely the $2$-uniform frames. The proof of Lemma 3.3 shows that any $3$-homogeneous ETF must be $3$-uniform. ETFs which have purely imaginary triple products must have $n=2d$ [ET00, Str08] and always exist when $2d\times 2d$ skew-symmetric conference matrices exist [DGS91]. However, we will see in Theorem 3.5 that only certain skew-symmetric conference matrices will lead to $3$-homogeneous ETFs.
As we will see in Theorem 3.5, $k$-homogeneity of ETFs for $k>1$ only holds in a few interesting cases. For $\geq 4, kk$-homogeneity basically only holds for orthonormal bases and simplices or because (nk)(n-k)-homogeneity implies $k$-homogeneity. The remaining cases to study are $2$-homogeneous ETFs which are not $2$-transitive (otherwise one could look to [IM22, DK23, IM24]) and $2$-transitive ETFs which are $3$-homogeneous. The latter case is completely characterized below. The $(2,4)$-SIC is the unique (up to switching equivalence and vector permutation, cf. [Zhu10]) ETF of $4$ vectors in $\mathbb{C}^{2}$ and the real $(3,6)$-ETF is the unique (up to switching equivalence and vector permutation, cf. [BK06]) real $(3,6)$-ETF.
Theorem 3.5.
The only ETFs with $3$-homogeneous line symmetry groups are the $(2,4)$-SIC, the $(d,2d)$-ETFs with $d=q+1$ where $q\equiv 3\bmod 4$ is a prime power arising from Paley skew-conference matrices, orthonormal bases, and simplices, which are all $2$-transitive. The only ETFs with $4$-homogeneous line symmetry groups are the $(2,4)$-SIC, the real $(3,6)$-ETF, orthonormal bases, and simplices.
Proof.
We consider $(d,n)$-ETFs, using [FM16] as a reference for $(d,n)$ which yield ETFs. By [LW65], if $n\geq 2k$, then any $k$-homogeneous group action on a set of size $n$ is (k1)(k-1)-transitive. Thus, all $3$-homogeneous ETFs (lines or vectors) with $n\geq 6$ are also $2$-transitive. By Lemma 3.4, the only ETFs which are not simplices or orthonormal bases but have $3$-homogeneous line symmetry group must have $n=2d$. It follows from [IM24] that the only $2$-transitive ETFs with $n=2d\geq 6$ and purely imaginary inner products have $d=q+1$ with $q\equiv 3\bmod 4$ and are generated by Paley skew-conference matrices. These have line symmetry group $\operatorname{PSL}(2,q)$ acting necessarily faithfully and transitively, where (see, e.g., [Kan72]) $\operatorname{PSL}(2,q)$ acts $3$-homogeneously. The only ETF with $n<6$ which is not an orthonormal basis or simplex is the $(2,4)$-SIC, which is $1$-homogeneous as a SIC and thus 3=(41)3=(4-1)-homogeneous as well. Note that the $(2,4)$-SIC has $2$-transitive line symmetry group [Zhu15]; thus, all of the ETFs with $3$-homogeneous line symmetry group are also $2$-transitive.
Applying [LW65] again, we see that all $4$-homogeneous ETFs (lines or vectors) with $n\geq 8$ are also $3$-transitive, which means they are simplex ETFs or orthonormal bases (Lemma 3.2). Similarly, ETFs with $4\leq n<8$ which are orthonormal bases or simplex ETFs trivially have $4$-homogeneous line symmetry groups (Lemma 3.1). The remaining ETFs which have $n<8$ are the $(2,4)$-SIC, the $(3,6)$-ETFs, the $(3,7)$-ETF(s) and $(4,7)$-ETF(s). The $(2,4)$ SIC is trivially $4$-homogeneous since the identity maps the group of four lines to the group of four. If a (not necessarily real) $(3,6)$-ETF is $4$-homogeneous, then it is $2$-homogeneous. Since $6$ is not prime power, any $2$-homogeneous $(3,6)$-ETF must be $2$-transitive [Kan69b]. The real $(3,6)$-ETF corresponds to the antipodal vertices of the regular icosahedron and thus has line symmetry group isomorphic to $\operatorname{PSL}(2,5)$, which acts $2$ transitively on the lines (see, e.g., Theorem 1.3 [IM24] which also gives uniqueness up to equivalence of the $(3,6)$-ETFs with $2$-transitive symmetry), in turn implying that the automorphism group is $2$-homogeneous and thus 4=(62)4=(6-2)-homogeneous. The remaining ETF parameters are the $(3,7)$- and $(4,7)$-ETFs. We have already shown that neither is $3$-homogeneous; thus, neither can have a 4=(73)4=(7-3)-homogeneous line symmetry group. ∎
As with the characterization of $4$-homogeneity, in the remaining cases where $k>4$, any $k$-homogeneous ETF with $n\geq 2k$ will be (k1)(k-1)-transitive and thus can only be a simplex ETF or orthonormal basis. For $n<2k$, the ETF will be (nk)(n-k)-homogeneous with nk<kn-k<k. Thus, inductively, the only possibilities are $1$-homogeneous ETFs (i.e., the ETF lines are spun up as the orbit of a group with no additional symmetries), $2$-homogeneous ETFs, or any of the $3$ or $4$-homogeneous ETFs enumerated above.
There are skew-conference matrices which are not Paley (e.g., [CD07]). Thus, being an ETF with purely imaginary triple products is not sufficient to be $3$-homogeneous.
Corollary 3.6.
The only ETFs with $3$-homogeneous vector symmetry groups are orthonormal bases and the regular simplex ETFs. The only ETFs with $4$-homogeneous vector symmetry groups are the $(2,4)$-SIC, orthonormal bases, and simplices.
Proof.
By following the cases in the proof of Theorem 3.5, we see that the only possible ETF which has a $3$-homogeneous vector symmetry group which is not $2$-transitive is the $(2,4)$-SIC. Although any SIC is the orbit of a fiducial under a unitary system, the unitary system is a projective unitary representation and thus does not act transitively on the vectors. Hence, the vector symmetry group of the $(2,4)$-SIC is neither $1$-homogeneous nor 3=(41)3=(4-1)-homogeneous (Lemma 3.3). All other ETFs with ETFs with $3$-homogeneous line symmetry groups must be $2$-transitive, and hence Lemma 3.2 applies.
By Theorem 3.5, we just need to check the $(2,4)$-SIC and the $(3,6)$ real ETF for $4$-homogeneous automorphism vector groups. The $(2,4)$-SIC is trivially $4$-homogeneous. If the $(3,6)$ real ETF had a $4$-homogeneous vector symmetry group, then (Lemma 3.3) it would be 2=(64)2=(6-4)-homogeneous, which (Lemma 3.2) cannot happen since it is neither an orthonormal basis or a regular simple ETF. ∎
4 Dave Larson’s AMS Memoirs
The preceding results mirror several themes in [HL00, DL98]; namely, that group actions are a powerful tool to construct frames with desirable properties, that Naimark complementation can be used to simplify the search for frames with particular properties, and that associated groups other than those used to generate the frame are powerful tools in understanding the frame properties.
Many infinite frames of note—like Gabor/time-frequency and wavelet—are generated as the (subsampling) of the orbit of a (projective) unitary representation and were a main focus of Dave Larson’s two AMS Memoirs [HL00, DL98]. A Gabor frame with time-frequency shifts over a lattice would be an example of an infinite frame with a transitive automorphism group of lines. We consider for a moment automorphism groups of such frames over infinite dimensional (separable) Hilbert spaces. Rephrasing Theorem 3.8 in [HL00]: If a tight frame $\Phi$ is the orbit of a unitary representation of a group $H$, then the Naimark complement is also the orbit of some unitary representation of the same group. Such a group $H$ is a subgroup of the automorphism groups of the lines and the vectors; so, this result is related to the fact that Theorem 1.14 implies that the automorphism groups of a finite tight frame and a Naimark complement are the same. Also, Corollary 3.10 of [HL00] yields that if a tight frame with frame bound $1$ is the orbit of a single bounded unitary operator, then it is unitarily equivalent to $\{e^{2\pi in\cdot}|_{E}\,:\,n\in\mathbb{Z}\}$ for $E\subset\mathbb{T}$, which is related to submatrices of Fourier matrices always yielding tight frames.
Assume that $\Phi=(\varphi_{j})_{j\in J}$ is a tight frame with frame bound $A$ for an infinite dimensional separable Hilbert space $\mathcal{H}$. Then $J$ is countably infinite. If $\Phi$ were $2$-homogeneous, then $\left|\langle\varphi_{i},\varphi_{j}\rangle\right|=\left|\langle\varphi_{k},\varphi_{\ell}\rangle\right|$ for all $i\neq j$ and $k\neq\ell$. Call this number $a$. By the definition of a tight frame (Definition 1.1), we see that
$\left\|\varphi_{k}\right\|^{2}=\frac{1}{A}\sum_{j\in J}\left|\langle\varphi_{k},\varphi_{j}\rangle\right|^{2}=\frac{1}{A}\left(\left\|\varphi_{k}\right\|^{2}+\sum_{j\neq k}a^{2}\right).$
Since the left-hand-side is a finite number, the right-hand-side must converge, meaning that $a=0$. Thus, the only infinite $2$-homogeneous infinite frames are orthogonal bases.
An important concept introduced in [DL98] (with additional contemporary publications [DL98, HWW96, HWW97, FW96, DLS97, DLS98]) is wavelet sets. (Parseval) wavelet sets are measurable subsets of $\mathbb{R}^{d}$ for which the integer (lattice) translates satisfy one condition and the dyadic dilations satisfy another condition. When $X$ is a (Parseval) wavelet set, then the inverse Fourier transform of the indicator function of $X$ generates a (wavelet tight frame with frame bound $1$) a wavelet orthonormal basis. The simple geometric characterization of wavelet sets allowed the discovery of wavelets with surprising features. We may view the construction of Fourier ETFs as a kindred concept. By instead zeroing out rows rather than deleting them, we could restate Theorem 1.10 as meaning that modulating an indicator vector of a set yields a ETF for the span if and only if the set is a difference set in the group used to generate the Fourier matrix. As in Example 1.21, taking all of the shifts of any difference set (called the development) always yields a BIBD.
Finally, we note that one of the key tools developed in [DL98] was local (to a frame vector) commutants. This led to the analysis of the von Neumann algebra generated by the orbit of the left regular representation of a group $G$ being a tool to understand whether other representations of $G$ can yield frames [HL00]. A related idea is the use of Schurian association schemes to understand finite ETFs in [IM22, IM24]. It is possible that the similarities between these approaches could be leveraged to create further tools to characterize ETFs.
4.1 Problems
In honor of the “Problems” offered in Dave Larson’s Memoirs [HL00, DL98], we consider a handful of questions that this manuscript leads to.
Corollary 2.20 gives that the bender of any Paley ETF is a BIBD. However, it does not give information on the parameters of the BIBD. The bender of a Paley ETF $\Phi_{p}$ for $p$ a prime is simply the set of all subsets of size $(p+1)/2$. Calculations from Harley Meade show that the spark of $\Phi_{27}$ is $8$ and the bender is a $(27,8,28)$-BIBD. Calculations for Paley ETFs with larger $q$ were not tractable.
Question 4.1.
What are the spark and bender of the Paley ETFs $\Phi_{q}$ with non-cyclic additive groups?
Dangerously powered by one infinite class and one example outside of this class, we propose the following conjecture.
Conjecture 4.2.
Let $\Phi_{q}$ be a Paley ETF with $q=p^{s}$. Then the spark of $\Phi_{q}$ is $((p+1)/2)^{s}$ and the bender is a $(p^{s},((p+1)/2)^{s},\lambda)$-BIBD.
To aid in solving the conjecture, there is code [Kin] to allow one to generate Paley ETFs of arbitrary size. This code also generates the symmetries $T$ and $\Pi$. There is additionally a function to construct ETFs from any difference sets.
We note that the computation of the Gram matrix of $\Phi_{p}$ for $p$ prime in the proof of Lemma 2.11 shows that the $\Phi_{p}$ are neither of two other categories of ETFs which are in some sense $2$-transitive-adjacent, i.e., being DRACKN or roux [CGSZ16, IM22]. This leads to the following question.
Question 4.3.
Are there any ETFs which are simultaneously $2$-homogeneous and DRACKN or $2$-homogeneous and roux but not $2$-transitive?
It is a fact that any linear isometry of $\mathbb{C}^{d}$ with respect to $\ell^{p}$ with $p\neq 2$ is a generalized permutation matrix, i.e., a product of a diagonal unitary and a permutation matrix [LS94]. It is also a fact that any element $U$ of the automorphism group of a $(d,n)$-Fourier frame necessarily satisfies $\left\|U\varphi_{j}\right\|_{p}=\left\|\varphi_{j}\right\|_{p}$ for all $1\leq p\leq\infty$ with the $\varphi_{j}$ being a spanning set for $\mathbb{C}^{d}$. Being an isometry for $p\neq 2$ on a spanning set is not sufficient to prove that it extends to an isometry on the entire space. However, the computed automorphisms in this paper are all generalized automorphisms.
Question 4.4.
Are the elements of the automorphism group of a Fourier ETF always generalized permutations?
We point to code from Joey Iverson [Ive] to determine the vector and line symmetry groups of frames as a resource to help answer this question.
Acknowledgments
The author is especially grateful to Joey Iverson, for fruitful discussions and GAP code [Ive] to run some experiments. The author is also thankful to Harley Meade, who ran Joey’s code to help formulate hypotheses and explicitly proved that groups related to $\{M^{k}T^{\ell}\}$ act $2$-homogeneously on ETFs related to prime-order Paley ETFs, to Dustin Mixon for helpful feedback, and to the anonymous referees for their thoughtful suggestions which lead to great improvements in the text. Most of all, the author is grateful to Dave Larson for introducing her to frames twenty-something years ago.